### v3

# Jinhong Li1    Yiyang Geng1    Qiuping Wang1    Shujie Han2    Patrick P. C. Lee1

Abstract
Zoned Namespace (ZNS) defines a new abstraction for host software to flexibly manage storage in flash-based SSDs as append-only zones. It also provides a Zone Append primitive to further boost the write performance of ZNS SSDs by exploiting intra-zone parallelism. However, making Zone Append effective for reliable and scalable storage, in the form of a RAID array of multiple ZNS SSDs, is non-trivial, since Zone Append offloads address management to ZNS SSDs and requires hosts to specifically manage RAID stripes across multiple drives. We propose ZapRAID, a high-performance log-structured RAID system for ZNS SSDs by carefully exploiting Zone Append to achieve high write parallelism and lightweight stripe management. ZapRAID adopts a group-based data layout with a coarse-grained ordering across multiple groups of stripes, such that it can use small-size metadata for stripe management on a per-group basis under Zone Append. It further adopts hybrid data management to simultaneously achieve intra-zone and inter-zone parallelism through a careful combination of both Zone Write and Zone Append primitives. We implement ZapRAID as a user-space block device, and evaluate ZapRAID using microbenchmarks, trace-driven experiments, and real-application experiments. Our evaluation results show that ZapRAID achieves high write throughput and maintains high performance in normal reads, degraded reads, crash recovery, and full-drive recovery.
1 Introduction
Flash-based solid-state drives (SSDs) have been widely deployed as they have better performance and reliability than hard-disk drives (HDDs). Conventional SSDs expose a block interface, implemented inside the flash translation layer (FTL), to a host for accessing flash storage. However, such a block interface also introduces costly internal flash management operations (e.g., device-level garbage collection and address translation) that lead to unpredictable I/O performance [19] and reduced flash endurance [5]. The recently proposed NVMe Zoned Namespace (ZNS) interface [77, 5] abstracts flash-based SSDs as append-only zones, so that the FTL and its costly operations are eliminated, and the storage management is now shifted to the host. Compared with conventional SSDs, ZNS SSDs are shown to achieve higher write throughput, lower tail read latencies, and less device-level DRAM usage [5].
ZNS SSDs support two write primitives, namely Zone Write and Zone Append, where Zone Append provides opportunities for further performance gains by exploiting intra-zone parallelism. Specifically, a ZNS SSD tracks per-zone write pointers, such that any write to a zone must specify the same offset indicated by the write pointer. Zone Write requires the host to specify the block address when writing a block, as in conventional SSDs. This provides backward compatibility with existing applications. However, in order to match the offset of a Zone Write with the on-device write pointer of a zone, the host can only issue one request to each zone at a time, thereby limiting intra-zone parallelism. In contrast, Zone Append eliminates the need for the host to specify block addresses in writes by fully offloading address management to ZNS SSDs. In Zone Append, the host only specifies the zone to which a write is issued, and the ZNS SSD returns the address to the host upon write completion. Thus, the host can issue multiple Zone Append commands within one zone to exploit the intra-zone parallelism for improved write performance.
Despite the performance gains of Zone Append, its offloading of address management to ZNS SSDs implies that the host not only cannot directly specify the block addresses of writes in a zone, but it also cannot control the ordering of writes in concurrent Zone Append commands. This creates new challenges to storage system designs that require host-level address management, particularly when applying Redundant Array of Independent Disks (RAID) [52] to form an array of multiple ZNS SSDs for reliable and scalable storage. Specifically, in a traditional RAID array, a RAID controller runs atop multiple drives and organizes data in stripes, each of which encodes a set of data blocks into parity blocks and distributes the data and parity blocks across drives. For efficient data repair, the RAID controller statically assigns the same block address in each drive for the blocks of the same stripe, so that the repair of any lost block can directly retrieve the available data and parity blocks of the same stripe from other drives for decoding. However, under Zone Append, it is infeasible for the RAID controller to manage block addresses in a zone; instead, the RAID controller needs to maintain dedicated address mapping information to specify the block locations for each stripe, which unavoidably incurs performance penalties to stripe management.
We present ZapRAID, a high-performance RAID system for ZNS SSDs by carefully exploiting Zone Append to achieve high write performance via intra-zone parallelism; in the meantime, ZapRAID supports lightweight stripe management in terms of: (i) querying stripe metadata during degraded reads, crash recovery, and full-drive recovery, (ii) memory usage for indexing, and (iii) persistent storage for metadata. ZapRAID builds on log-structured RAID (Log-RAID) [11, 14, 23, 30], which issues sequential writes to a RAID array, and extends Log-RAID with a novel group-based data layout. Specifically, ZapRAID partitions stripes into stripe groups and issues Zone Append to the stripes within the same stripe group for high write performance. The group-based data layout organizes stripes with a coarse-grained ordering and enables ZapRAID to manage stripes efficiently on a per-group basis.
From our measurement study (§2.2), we observe that Zone Append effectively exploits intra-zone parallelism when a workload is dominated by small writes and when writes are issued to a single zone of a ZNS SSD at a time, while Zone Write is beneficial when a workload is dominated by large writes or when writes are concurrently issued to multiple zones of a ZNS SSD (in which Zone Write exploits better inter-zone parallelism for large writes). To this end, ZapRAID adopts hybrid data management to manage small and large writes across multiple zones in a ZNS SSD, so as to simultaneously achieve intra-zone and inter-zone parallelism through a careful combination of both Zone Append and Zone Write primitives.
In summary, we make the following contributions:
  • We conduct a measurement study to carefully examine the trade-off between Zone Append and Zone Write on a ZNS SSD (§2.2).
  • We design ZapRAID, a high-performance Log-RAID system for ZNS SSDs. ZapRAID builds on two core ideas: (i) a group-based data layout under Zone Append (§3.2) and (ii) hybrid data management (§3.3) under a combination of Zone Write and Zone Append. ZapRAID further addresses reliability in terms of metadata persistence and crash consistency (§3.4).
  • We prototype ZapRAID as a user-space block device with the Storage Performance Development Kit (SPDK) [65], such that ZapRAID exports a block interface for compatibility with general applications and can parallelize its internal operations with multiple SPDK threads.
  • We evaluate our ZapRAID prototype on real ZNS SSD devices using microbenchmarks, trace-driven experiments, and real-application experiments. When writes are issued to a single zone per ZNS SSD, ZapRAID increases the write throughput by up to 77.2% and reduces the tail latency by up to 36.4% compared with the exclusive use of Zone Write, and also scales to high write throughput under FEMU emulation [39]. When writes are issued to multiple zones per ZNS SSD, ZapRAID always maintains the highest throughput and lowest tail latency compared with the exclusive use of either Zone Write or Zone Append. Furthermore, ZapRAID maintains high performance in normal reads, degraded reads, crash recovery, and full-drive recovery.
We now make the source code available on https://github.com/adslabcuhk/zapraid.
2 Background and Motivation
We provide the basics of ZNS (§2.1) and study the trade-off between Zone Write and Zone Append (§2.2). We also review the background of Log-RAID and state the challenges of deploying Log-RAID on ZNS SSDs (§2.3).
2.1 Zoned Namespace (ZNS)
Figure 1: Architecture of a ZNS SSD based on the Western Digital Ultrastar DC ZN540 model [70].
The Zoned Namespace (ZNS) interface [77, 5] is a new NVMe Command Set [50] that exposes flash-based SSD internals to the host for host-level storage management. Figure 1 shows the architecture of a ZNS SSD based on the Western Digital Ultrastar DC ZN540 model, which we use in this paper and also in previous studies (e.g., [4, 31, 35, 60, 16]). The ZNS interface abstracts a ZNS SSD as append-only zones, each of which has a maximum size called the zone capacity. Each zone organizes data in logical blocks (or blocks in short in this paper), and each block is mapped to a physical flash page. Each flash page also has an out-of-band area for metadata storage (of size tens of bytes) located alongside the page content. Note that the out-of-band area is also available in general NVMe SSDs, and its space is not included in the total zone capacity. In this paper, we assume that the block size is 4 KiB and the out-of-band area size is 64 bytes, both of which can be configured in the ZN540 model when a ZNS SSD is formatted.
Within a zone, the blocks can be randomly read, but must be sequentially written; meanwhile, multiple zones can be read/written concurrently. Specifically, a ZNS SSD maintains a per-zone write pointer that specifies the offset of the next write. It mandates that any new write to a zone must be issued to the offset referenced by the write pointer. Upon the completion of each write request, the write pointer is incremented by the number of written blocks.
Each zone in a ZNS SSD is associated with a state. An empty zone means that the zone has not been written any block. Once a block is written to an empty zone, the zone becomes an open zone and can accept new writes before it is full. Typically, the number of open zones in a ZNS SSD is bounded. Once the number of blocks written to a zone reaches the zone capacity, the zone becomes a full zone. The host can also explicitly finish an open zone to turn it into a full zone. To overwrite the data in a zone, the host must first explicitly reset the zone to erase any data in a block and rewind the write pointer to the beginning of the zone, and the zone becomes an empty zone after reset.
2.2 Zone Write versus Zone Append
Recall from §1 that ZNS provides the Zone Write and Zone Append primitives for a host to issue writes to a zone. Their key difference is that Zone Write requires the host to issue writes serially and specify the offset in each write, while Zone Append offloads address assignment to ZNS SSDs and enables the host to issue concurrent writes that exploit intra-zone parallelism. Existing ZNS devices and host software typically allow only one outstanding Zone Write per zone in order to maintain the write pointer invariant; otherwise, concurrent Zone Write and Zone Append commands may violate the sequential-write requirement.
We conduct evaluation on a single ZN540 ZNS SSD [70] in our testbed (see §5.1 for testbed details) to show how Zone Append improves write performance over Zone Write. We issue write requests of 4 KiB, 8 KiB, and 16 KiB to the ZNS SSD. We also vary the number of open zones in the ZNS SSD. We write a total of 512 GiB of data and measure the write throughput. We report the average throughput results over five runs, with the error bars showing the 95% confidence intervals under the Student’s t-distribution (note that some error bars may be invisible due to small deviations). Figure 2 shows the write throughput results.
Refer to caption Refer to caption Refer to caption
(a) 4 KiB requests (b) 8 KiB requests (c) 16 KiB requests
Figure 2: Write throughput of Zone Write and Zone Append versus the number of open zones under various request sizes.
Writes to a single open zone. We first consider the case when the number of open zones is one. In this case, we issue writes to a single zone at a time; if the zone is full, we issue writes to a different zone. Zone Write achieves a write throughput of 337.6 MiB/s, 613.6 MiB/s, and 1,050.0 MiB/s for 4 KiB, 8 KiB, and 16 KiB requests, respectively. Recall that a ZNS SSD can have only one outstanding Zone Write at any time. We find that the write throughput of Zone Write cannot further increase even if we issue multiple concurrent write requests of Zone Write. In contrast, Zone Append achieves a write throughput of 541.5 MiB/s, 1,026.6 MiB/s, and 1,050.1 MiB/s for 4 KiB, 8 KiB, and 16 KiB requests, respectively, with four concurrent write requests of Zone Append [^1] The ZN540 specification [70] states that the write throughput can reach 2,000 MiB/s. This number corresponds to the best-case performance and is obtained when large requests (128 KiB) are issued to all open zones.. This shows that Zone Append can increase the write throughput via intra-zone parallelism.
Note that in this single-zone experiment, even if we issue more than four concurrent write requests, the write throughput of Zone Append for 4 KiB and 8 KiB requests does not further increase, as it already saturates intra-zone parallelism for this specific ZN540 model. Also, Zone Write already achieves the maximum write throughput of a zone in ZN540 for 16 KiB requests, so Zone Append does not see more growth.
Writes to multiple open zones. We further consider the case when the number of open zones is larger than one, so as to examine how inter-zone parallelism affects the effectiveness of Zone Write and Zone Append. We create multiple open zones and issue writes to the open zones in parallel. We set the number of concurrent write requests for each open zone as one and four for Zone Write and Zone Append, respectively, based on the previous results in the single-zone experiment. From Figure 2, we observe that the write throughput of Zone Write is generally higher when there are more open zones (e.g., its write throughput reaches 777.1 MiB/s under six open zones for 4 KiB write requests (Figure 2(a))). On the other hand, Zone Append can only achieve a maximum write throughput of 577.5 MiB/s under two open zones for 4 KiB requests (Figure 2(a)). A possible reason is that the current firmware implementation of Zone Append is more computationally intensive when writes are issued to more open zones in the ZN540 model, so Zone Append has even lower write throughput than Zone Write under a larger number of open zones due to the limited computational power in existing hardware. We conjecture that this limitation could be addressed in future ZNS SSD models. When the request size is 8 KiB (Figure 2(b)), Zone Write can scale to 1,430.7 MiB/s under eight open zones, while Zone Append can only achieve a maximum write throughput of 1,058.6 MiB/s under two open zones. When the request size is 16 KiB (Figure 2(c)), Zone Write and Zone Append have very small performance differences (e.g., up to around 1,750 MiB/s under eight open zones), implying large writes can effectively exploit inter-zone parallelism in both Zone Write and Zone Append.
Main observations. We observe that for small writes (e.g., of size 4 KiB or 8 KiB), Zone Append achieves higher write throughput than Zone Write in the single-zone setting by exploiting intra-zone parallelism, while Zone Write scales better and achieves higher write throughput than Zone Append when the number of open zones increases by better exploiting inter-zone parallelism. For large writes (e.g., of size 16 KiB), both Zone Write and Zone Append have similar write throughput for different numbers of open zones. A RAID system based on ZNS SSDs should adequately combine Zone Append and Zone Write to accommodate the scenarios with different numbers of open zones. Even though Zone Append only shows high performance when write requests are issued to one or a few open zones, it is still beneficial for the following scenarios. First, shared storage systems need performance isolation [9, 20, 27]. Specifically, since a zone is typically mapped to only a subset of physical flash channels or chips [3], a shared storage system based on ZNS SSDs should allocate only a limited number of open zones to applications that do not have high write bandwidth demands. This isolates the write traffic of the applications to a subset of flash channels and limits their impact on other co-located applications. Second, it is common to have mixed request sizes in many real-world workloads. For example, practical cloud block storage workloads have 60% of write requests that are at most 4 KiB and 25% of write requests that are at least 16 KiB [41]. By separating small and large writes into different open zones (§3.3), we can use Zone Append and Zone Write for different groups of writes to improve the overall performance (§5.4). Thus, Zone Append can help such applications achieve high write performance through intra-zone parallelism and sustain the bursts of a large number of write requests if necessary, even with only one or a few open zones per ZNS SSD being used.
2.3 Log-structured RAID (Log-RAID)
RAID [52] forms an array of storage drives to improve capacity, performance, and fault tolerance in storage. We focus on parity-based RAID, which organizes the RAID storage space in units of stripes. Each stripe contains multiple chunks that span across all drives (one chunk per drive), where each chunk contains a fixed number of contiguous blocks. A $(k+m)$-RAID array forms a stripe with $k+m$ chunks by encoding a set of $k$ data chunks into $m$ parity chunks, such that any $k$ out of the $k+m$ chunks of the same stripe can decode the $k$ data chunks. It distributes the $k+m$ chunks of each stripe across $k+m$ drives, so as to provide fault tolerance against any $m$ failed drives. To make data repair efficient, RAID performs static mapping, which aligns each data or parity chunk of the same stripe at the same offset of a drive. Thus, when recovering any lost chunk, the RAID controller can deterministically retrieve $k$ available chunks of the same stripe at the same offset from other drives for decoding.
Traditional RAID is designed for HDDs [52], which adopt in-place updates to overwrite blocks. In contrast, SSDs adopt out-of-place updates. Small random writes to SSDs are shown to trigger frequent device-level garbage collection that degrades both I/O performance and flash endurance [10, 28, 48]. Parity-based RAID is even more vulnerable to small random writes, which trigger parity updates that further aggravate performance and endurance degradations. Thus, some studies [11, 14, 23, 30] propose Log-RAID, which applies the log-structured file system design [59] to SSD RAID by issuing sequential host-level writes.
Log-RAID manages stripes in append-only segments. Each segment holds a number of stripes (up to some pre-specified capacity) and is mapped to $k+m$ fixed-size contiguous areas that reside in $k+m$ drives (one contiguous area per drive), so as to provide fault tolerance against any $m$ failed drives. A segment is open if it does not reach its capacity and hence can accept new writes. Log-RAID aggregates newly written chunks as new stripes. It writes the new $k+m$ chunks of each stripe to the same offset of the $k+m$ contiguous areas in an append-only manner, such that $k+m$ chunks of the same stripe are aligned at the same offset of the $k+m$ contiguous areas (i.e., static mapping). If an open segment reaches its full capacity, Log-RAID seals it into a sealed segment, and creates a new open segment from the free contiguous areas in the underlying drives.
By writing blocks as new stripes in an append-only manner, Log-RAID needs to perform garbage collection to reclaim the space from stale blocks. When garbage collection is triggered (say, when the available space drops below some threshold), Log-RAID selects a sealed segment by some policy (e.g., using a greedy algorithm to select the one with the largest fraction of stale blocks), rewrites all non-stale blocks into a new open segment, and releases the space of the selected sealed segment for reuse.
Unlike traditional RAID, Log-RAID necessitates dedicated in-memory data structures for stripe management. Specifically, it keeps a segment table to map each segment to its corresponding $k+m$ contiguous areas. It also keeps an address table to map each block address (identified by the application) to the corresponding segment, drive, and offset within the contiguous areas. To repair a lost block, Log-RAID first identifies the segment and drive in which the lost block resides from the address table. It then identifies the contiguous areas of the other drives for the segment from the segment table. It retrieves $k$ available chunks at the same offset in other contiguous areas to decode the lost chunk (as in the static mapping in traditional RAID), and finally recovers the lost block.
Log-RAID is a natural fit for ZNS SSDs by mapping each contiguous area of a drive to a zone, which organizes writes in an append-only manner. However, applying Zone Append in Log-RAID is non-trivial. Since Zone Append offloads address management to ZNS SSDs, the chunks of the same stripe may now be written to different offsets within the zones. To repair any lost block, we can no longer directly retrieve the available blocks from the same offset of other zones based on static mapping; instead, we need to keep track of the offsets of the chunks of the same stripe. How to efficiently manage chunk locations when Zone Append is used in Log-RAID for ZNS SSDs is an important design issue.
3 ZapRAID Design
ZapRAID is an extended Log-RAID design for ZNS SSD arrays with three key design goals: (i) high performance, i.e., ZapRAID achieves high throughput and low latency via a combination of Zone Append and Zone Write; (ii) lightweight stripe management, i.e., ZapRAID keeps both low memory usage for index structures and low storage overhead for metadata persistence, while supporting fast metadata queries for locating the chunks of a stripe; and (iii) reliability, i.e., ZapRAID protects data and metadata against failures.
ZapRAID achieves the above goals with two main designs. The first design is a group-based data layout. Specifically, when ZapRAID uses Zone Append for intra-zone parallelism, it maintains a coarse-grained ordering of stripes by organizing stripes in stripe groups, such that it issues Zone Append for the stripes on a per-group basis, so as to efficiently manage stripes for each stripe group instead of in a system-wide manner. It further enforces reliability by maintaining consistency for stripes and index structures during crash recovery. The second design is hybrid data management. ZapRAID serves the write requests of small and large sizes using different segments, such that a segment is configured with a small (resp. large) chunk size to serve small (resp. large) writes. This allows ZapRAID to serve small writes with low latencies and maintain high overall write throughput. Also, by managing segments in a hybrid manner, ZapRAID can further apply Zone Append and Zone Write for different segments to achieve intra-zone and inter-zone parallelism, respectively.
3.1 Design Overview
Architecture. We first introduce the architecture of ZapRAID and show how it augments Log-RAID on an array of multiple ZNS SSDs, as shown in Figure 3. ZapRAID exposes a block-level volume that supports random reads/writes. We assume that the block size is 4 KiB. A user application can issue reads/writes of an arbitrary number of blocks, each of which is identified by a logical block address (LBA). Each drive corresponds to a ZNS SSD, and ZapRAID maps each contiguous area of a drive under Log-RAID to a zone, which only allows sequential writes. Let $Z$ be the total number of zones in a drive, so there are $Z$ segments in a ZNS SSD array. For example, a 4 TiB ZN540 ZNS SSD [70] configures $Z=$ 3,690 zones. Each of the drives, segments, zones, and stripes is associated with an identifier (ID).
Figure 3: ZapRAID architecture. Here, we show a (3+1)-RAID-5 array and the layout of a segment of four zones, where the parity chunks are rotated across the drives.
Segment organization. Each segment corresponds to $k+m$ zones in $k+m$ drives. In ZapRAID, each segment comprises three regions that span across $k+m$ drives: the header region and the footer region for keeping metadata for crash consistency (§3.4), and the data region for storing data and parity chunks. Specifically, the header region stores the RAID scheme (including RAID-5/RAID-6, $k$, and $m$) and the segment-specific information (including the zone IDs of all zones from the $k+m$ drives in the segment, the chunk size, and the stripe group size (defined in §3.2)). The footer region keeps the block metadata, including the LBA, write timestamp, and stripe ID of each block in the data region.
All three regions have pre-specified sizes. The header region contains exactly one stripe of $k+m$ chunks, each of which resides at the beginning of a zone. The data region contains a fixed number of stripes, denoted by $S$, following the header region. For each segment, each chunk contains a fixed number of blocks, denoted by $C$, so the number of blocks in the data region is $S\cdot C$. Suppose that the LBA size is 8 bytes, the write timestamp size is 8 bytes, the stripe ID size is 4 bytes, and the block size is 4 KiB. Thus, each block in the footer region can store the block metadata of $\lfloor\tfrac{4096}{20}\rfloor=204$ blocks, so the footer region occupies $\lceil\tfrac{S\cdot C}{204}\rceil$ blocks following the data region. For example, the capacity of a zone in a ZN540 ZNS SSD [70] is 1,077 MiB (or equivalently, 275,712 blocks). If the chunk size is 4 KiB (i.e., $C=1$), then the header, data, and footer regions occupy 1 block, 274,366 blocks, and 1,345 blocks in a zone, respectively.
When ZapRAID creates a new open segment across $k+m$ drives, it first writes the header region. It then writes the stripes to the data region. Depending on the RAID scheme, the parity chunks of all stripes may reside in a fixed drive (for RAID-4) or be rotated across the drives (for RAID-5 and RAID-6) (e.g., see the RAID-5 example in Figure 3). When the data region reaches its pre-specified size, ZapRAID writes the block metadata of all blocks in the segment to the footer region. Finally, it seals the segment.
Block metadata. ZapRAID stores the block metadata for each block in the out-of-band area of the corresponding flash page (§2.1) for persistence. Each block in a data chunk has its LBA, write timestamp, and stripe ID as its block metadata. ZapRAID provides fault tolerance for the block metadata: for LBAs and write timestamps, ZapRAID generates parity-based redundancy for them from all blocks of the data chunks in the same stripe and stores the parity results in the block metadata of the blocks of the parity chunks, while for stripe IDs, ZapRAID replicates them into all the data and parity chunks in the same stripe.
Recall that the footer region also stores the block metadata (i.e., the LBA, write timestamp, and stripe ID) for all blocks in the segment, so ZapRAID keeps two copies of block metadata (i.e., in the out-of-band area of each flash page and the footer region). We argue that both copies are necessary and serve different purposes: the block metadata in the out-of-band area associated with each block provides persistence for block writes, while the block metadata in the footer region provides fast crash recovery (§3.4).
In-memory items. ZapRAID keeps a number of in-memory in-flight stripes for newly written blocks. Each in-flight stripe is kept in memory until all of its $k$ data chunks and $m$ parity chunks are formed and persisted. To maintain durability, ZapRAID acknowledges the writes of an in-flight stripe only after the whole in-flight stripe is persisted (note that acknowledging the write of each block can lead to data loss if a drive storing an acknowledged block fails, but the parity chunks are not yet generated).
ZapRAID also maintains three in-memory index structures: (i) the segment table, which maps each segment ID to its corresponding $k+m$ zones (identified by the zone IDs in the respective $k+m$ drives), segment state, chunk size, and stripe group size; (ii) the logical-to-physical (L2P) table, which maps the LBA of each block issued by an application to the physical block address (PBA) (i.e., the segment ID, the drive ID, and the offset in the respective zone); and (iii) the compact stripe table, which maps each segment ID to the stripe IDs of all chunks in the segment. Both the segment table and L2P table are similarly found in Log-RAID. The compact stripe table is newly introduced to ZapRAID since Zone Append can make the chunks of the same stripe reside at different offsets across the drives (§2.3). We show how the compact stripe table is designed and how we reduce its size via the group-based data layout in §3.2. Note that ZapRAID ensures fault tolerance for the index structures by persisting the segment-to-zones mappings in the segment table into the header region of each segment and persisting the LBAs, write timestamps, and stripe IDs as block metadata into the out-of-band area associated with each block.
Offloading L2P table entries to ZNS SSDs. The L2P table incurs large memory footprints if it keeps all entries in memory. For example, our current implementation supports a RAID array of four 4 TiB ZN540 ZNS SSDs (i.e., a total capacity of 16 TiB) and uses a 4-byte L2P entry for each 4 KiB block. Thus, the total memory size of the L2P table is 16 GiB. Such large L2P table footprints are common in all Log-RAID designs [11, 14, 23, 30]. Some Log-RAID designs keep part of the L2P table entries on disk to mitigate the memory overhead [11, 14].
ZapRAID by default keeps the whole L2P table in memory, yet it also supports the offloading of part of the L2P table entries to ZNS SSDs. It uses the CLOCK algorithm [15] to evict the non-recently used L2P entries to SSDs. Specifically, it divides the L2P table entries into multiple groups, each of which has 1,024 contiguous entries that can be fit into a 4 KiB block (assuming that each entry is of size four bytes). It maintains an in-memory bitmap, in which each bit tracks if the corresponding entry group has been accessed, and a pointer that references the entry group to be considered. It sets the tracked bit to one if any entry in the corresponding entry group has been accessed. Suppose that the L2P table size reaches the memory size limit. ZapRAID then searches for an entry group to evict, starting from the one referenced by the pointer. If the tracked bit of an entry group is one, ZapRAID resets the tracked bit to zero, moves the pointer to the next entry group, and continues the search; otherwise, if the tracked bit of the entry group is zero, ZapRAID selects the entry group to evict, moves the pointer to the next entry group, and stops the search.
ZapRAID maps the entry group into a 4 KiB mapping block. It writes the mapping block together with other user-written blocks, so as to avoid consuming additional open zone resources. It maintains a separate in-memory mapping table to record the PBA of each on-SSD mapping block. Note that the size of the in-memory mapping table is much smaller, as each of its entries can track 1,024 L2P table entries (e.g., in our case, the size of the in-memory mapping table is 16 MiB for a RAID array of four 4 TiB ZN540 ZNS SSDs). It also keeps the LBA of the first entry of the mapping block in the LBA field of the block metadata for crash recovery (§3.4).
It is important to distinguish between the mapping blocks and user-written blocks during the crash recovery of the L2P table (§3.4). Thus, ZapRAID marks the least significant bit of the LBA field in the block metadata as one for a mapping block. It is feasible since the LBA field of a user-written block is currently aligned at multiples of 4 KiB, and its 12 least significant bits are all zeros.
3.2 Group-Based Data Layout
ZapRAID exploits Zone Append for intra-zone parallelism, yet it needs to track the offsets of the chunks of the same stripe under Zone Append. Thus, ZapRAID adopts a group-based data layout to organize the stripes of a segment with a coarse-grained ordering for low stripe management overhead. Specifically, for a segment where Zone Append is issued, ZapRAID partitions a fixed number of contiguous stripes, denoted by $G$, within the segment into stripe groups, where $G$ is a configurable parameter. For each stripe group, it first issues Zone Append for all stripes within the same stripe group, such that all chunks of each stripe must reside in the same stripe group but may reside in different offsets within the stripe group. Each stripe group is in the same offset ranges across all zones, so the offset ranges of its blocks can be identified via static mapping. Most importantly, ZapRAID only needs to track a small number of stripes within each stripe group, so it can use fewer bits for the metadata information to save significant memory usage. Note that we treat $G=1$ as a special case in which each stripe group issues Zone Write instead of Zone Append for all stripes.
Figure 4 depicts one segment with $G=4$ stripes per stripe group in a (3+1)-RAID-5 array. Within the segment, the data region comprises a fixed number of stripe groups, each of which further comprises a fixed number of stripes. In general, the number of stripe groups in a segment is determined by both $S$ (i.e., the number of stripes in the segment) and $G$. Each stripe in a stripe group is associated with a unique stripe ID, which can be viewed as a sequence number of when the stripe is generated. Due to Zone Append, the chunks of the same stripe may reside in different offsets, as shown in Figure 4.
Figure 4: Group-based data layout of a (3+1)-RAID-5 array, where $G=4$. Each chunk is labeled with its stripe ID in the group, while the chunks with the same stripe ID in the same stripe group form a stripe.
ZapRAID tracks the stripe IDs of all stripe groups in the compact stripe table. Specifically, for each segment where Zone Append is issued, the compact stripe table stores a two-dimensional $(k+m)\times S$ matrix, in which each entry stores the stripe ID of each chunk in the segment. In the matrix, the $i$-th row corresponds to the zone of the $i$-th drive of the RAID array, and the $j$-th column corresponds to the $j$-th chunk in the segment. Given the LBA of a block, ZapRAID identifies the PBA from the L2P table, calculates the physical location of the chunk based on the PBA, and retrieves the stripe ID from the compact stripe table.
Trade-off analysis. The choice of $G$ determines the trade-off between the degree of intra-zone parallelism via Zone Append and the stripe management overhead. A larger $G$ allows more stripes to be issued via Zone Append in parallel, but it also increases the stripe management overhead. In the extremes, when $G=1$, all stripes are issued via Zone Write; when $G=S$, there is only one stripe group in each segment (i.e., the group-based data layout is disabled for Zone Append).
[…]
In contrast, the case of $G=S$ (i.e., without the group-based data layout) translates to 19 bits per stripe ID and hence 8.95 GiB of memory for the compact stripe table. Also, its high query overhead significantly increases the degraded read latency (§5.4).
Our prototype rounds up the length of a stripe ID to the nearest byte and represents a stripe ID in $\lceil\lceil\log_{2}G\rceil/8\rceil$ bytes. Representing stripe IDs in units of bytes, instead of bits, simplifies data alignment without compromising our conclusion. Thus, for $G=S$, our prototype consumes 3 bytes per stripe ID, leading to 11.3 GiB of memory for the compact stripe table.
3.3 Hybrid Data Management
If an application can write data to multiple open zones in a ZNS SSD, ZapRAID can create multiple open segments across the open zones and combine Zone Append and Zone Write to achieve both intra-zone and inter-zone parallelism. Recall that Zone Write achieves higher write throughput than Zone Append under multiple open zones (i.e., higher inter-zone parallelism) and maintains high write throughput for large writes. Unlike Zone Append, Zone Write can maintain static mappings in Log-RAID and hence does not require the compact stripe table for stripe management. Thus, the core idea of ZapRAID is to reserve an open segment for small writes to exploit intra-zone parallelism based on Zone Append, while favoring Zone Write for large writes to exploit inter-zone parallelism.
To this end, ZapRAID adopts hybrid data management. It classifies the open segments into two types, namely small-chunk segments and large-chunk segments, to serve small and large writes, respectively. It applies Zone Append to one of the small-chunk segments based on the group-based data layout (§3.2), and applies Zone Write to the remaining small-chunk segments and all large-chunk segments.
Segment management. Figure 5 shows how ZapRAID manages multiple open small-chunk and large-chunk segments. Let $N_{s}$ and $N_{l}$ be the numbers of open small-chunk segments and open large-chunk segments, respectively (i.e., the total number of open zones in each ZNS SSD is $N_{s}+N_{l}$). Also, let $C_{s}$ and $C_{l}$ be the sizes of a small chunk and a large chunk, respectively. All of $N_{s}$, $N_{l}$, $C_{s}$, and $C_{l}$ are configurable parameters. For example, we can set $C_{s}=8$ KiB for the small chunk size and $C_{l}=16$ KiB for the large chunk size. We can also set $N_{s}>1$ and $N_{l}>1$ to accommodate the real-world workloads that are mixed with small and large writes (e.g., in cloud block storage [41]).
Figure 5: Hybrid data management of a (3+1)-RAID-5 array with multiple open segments.
To determine if a write request is issued to a small-chunk or large-chunk segment, ZapRAID examines whether the size of the write request is smaller than a write request size threshold, which we now set as $C_{l}$. If so, ZapRAID assigns the write request to a small-chunk segment; otherwise, it assigns the write request to a large-chunk segment. Also, during garbage collection (§4), the rewritten data is written to a large-chunk segment to mitigate I/O overhead.
Choosing segments from in-flight stripes. ZapRAID maintains in-flight stripes separately for small and large chunks. It assigns new written blocks to an in-flight stripe for either small or large chunks based on the write request size (see above). When an in-flight stripe is formed, ZapRAID forwards the in-flight stripe to an open segment. Specifically, if the in-flight stripe holds large chunks, ZapRAID selects an open large-chunk segment (out of the $N_{l}$) in a round-robin manner and writes the in-flight stripe to the selected segment via Zone Write. On the other hand, if the in-flight stripe holds small chunks, ZapRAID further considers two cases of $N_{s}$. If $N_{s}=1$, it writes the in-flight stripe to the only open small-chunk segment via Zone Append. If $N_{s}>1$, it selects an idle small-chunk segment (out of the $N_{s}-1$) in a round-robin manner and writes the in-flight stripe to the selected segment via Zone Write; however, if there is no idle small-chunk segment, ZapRAID selects the only small-chunk segment reserved for Zone Append and writes the in-flight stripe to the segment via Zone Append.
3.4 Crash Consistency
We show how ZapRAID maintains consistency after a system crash, in the order of recovering the consistent states of the segment table, stripes, L2P table, and compact stripe table (for Zone Append only).
Recovery of the segment table. ZapRAID first identifies the open and sealed segments in the system, so as to recover the segment table. Specifically, for each ZNS SSD, ZapRAID checks the states of all zones. If a zone is open or full, ZapRAID retrieves its first block, which belongs to the header region of a segment (§3.1). ZapRAID then reconstructs the segment table with the information of all possible segments based on the first blocks of all open and full zones.
ZapRAID further checks the validity of each segment. There are two possible cases:
  • Case 1: the write pointers of all the zones in the segment are equal to or exceed the header region size. In this case, the segment is a valid segment.
  • Case 2: the write pointers of some zones are zero. This happens when ZapRAID either has created the segment but has only issued writes to some zones, or has reclaimed the segment and has reset some zones. In either case, ZapRAID simply resets all zones and discards the segment.
For each valid segment, if the write pointers of all zones are equal to the zone capacity, it is a sealed segment; otherwise, it is an open segment. In particular, if the data region of an open segment reaches its maximum size but the segment is not yet sealed, ZapRAID will retrieve all blocks from the data region to reconstruct the block metadata. It then writes the footer region and seals the segment.
Recovery of stripes. ZapRAID needs to recover the consistent states of all stripes by ensuring that each stripe has persisted $k$ data chunks and $m$ parity chunks. Specifically, ZapRAID examines whether each in-flight stripe has all its $k+m$ chunks persisted. Under the group-based data layout (note that $G=1$ means the use of Zone Write), ZapRAID ensures that only the stripes in the latest stripe group of an open segment can be the in-flight stripes. It retrieves all the chunks in the latest stripe group and examines the stripe IDs in their block metadata. ZapRAID discards all partially persisted stripes with fewer than $k+m$ persisted chunks. Recall that ZapRAID acknowledges the writes of an in-flight stripe only after persisting all $k+m$ chunks of the stripe (§3.1), so discarding the partially persisted stripes does not incur data loss. Since ZNS SSDs cannot perform in-place writes to replace the partially persisted chunks, ZapRAID rewrites all fully persisted stripes in the segment to a new segment and reclaims the old segment.
Recovery of the L2P table and the compact stripe table. After ensuring the consistency of all stripes, ZapRAID recovers the L2P table and the compact stripe table (the latter is for segments that use Zone Append). Recall that for each sealed segment, ZapRAID stores the block metadata (i.e., the LBA, write timestamp, and stripe ID) of all blocks in the segment in the footer region, while for each open segment, ZapRAID also stores the block metadata in the out-of-band area associated with each block. To recover the L2P table and the compact stripe table, ZapRAID first retrieves the footer regions of all sealed segments and the data regions of all open segments. It examines all the block metadata and reconstructs the LBA-to-PBA mappings in the L2P table and the stripe IDs in the compact stripe table. For the L2P table, if multiple blocks have the same LBAs, ZapRAID keeps the one with the latest write timestamp.
To recover the L2P table when it is offloaded to ZNS SSDs (§3.1), ZapRAID extends the above recovery steps for the L2P table as follows. During crash recovery, ZapRAID examines the block metadata of each block and determines if the block is a user-written block or a mapping block (based on the least significant bit of the LBA). If the block is a user-written block, ZapRAID reconstructs the LBA-to-PBA mapping in the L2P table as described above; if the block is a mapping block, since the LBA in the block metadata refers to the LBA of its first L2P table entry, ZapRAID records the LBA-to-PBA mapping of the mapping block in a temporary in-memory table (if there exist multiple mapping blocks with the same LBA, only the one with the latest write timestamp is kept). ZapRAID then examines the reconstructed L2P table and removes any entry group that has a mapping block whose latest write timestamp is no smaller than that of any L2P table entry in the entry group (i.e., the mapping block has the latest L2P entries). Finally, ZapRAID reconstructs the in-memory bitmap and the in-memory mapping table for the on-SSD mapping blocks.
3.5 Complete Workflows
We explain the complete workflows of writes, normal reads, degraded reads, and full-drive recovery of ZapRAID.
[…]
ZapRAID needs to handle two specific cases during writes. First, if the written stripe is the first stripe in a stripe group, ZapRAID should ensure that all the stripes in the previous stripe group are completely persisted; if the stripe group is also the first one of a segment, the header region of the segment should be completely persisted. Second, if the stripe group is also the last one in the data region, ZapRAID writes the block metadata of all blocks in the segment into the footer region in the background. It also creates a new open segment and writes the segment information to the header region (without waiting for the completion of the writing of the footer region in the previous segment), so that the new open segment can serve new writes.
Reads. To read a block (identified by an LBA), ZapRAID queries the L2P table for its PBA (i.e., the segment ID, the drive ID, and the offset in the respective zone). Given the segment ID and drive ID, ZapRAID locates the zone from the segment table and retrieves the block from the specified offset of the zone.
Degraded reads. Suppose that ZapRAID performs a degraded read to a lost block. Based on the PBA of the requested block, ZapRAID checks whether the segment applies Zone Append or Zone Write. For Zone Append, ZapRAID queries the compact stripe table to find out the stripe ID of the requested block. Since all available chunks of the same stripe must reside in the same stripe group under the group-based data layout, ZapRAID searches for the offsets of the $k$ available chunks in the stripe group from the compact stripe table; note that the search is efficient due to the limited group size (§5.4). It then reads the available chunks from the other zones, decodes the lost chunk, and retrieves the requested block. For Zone Write, ZapRAID reads the available chunks from other zones at the same offset to recover the chunk and hence the requested block.
Full-drive recovery. When a drive fails, ZapRAID recovers the lost data into a new drive. It first identifies the segments that contain the lost zones in the failed drive by examining the segment-to-zones mappings in the header regions of all stored segments. For each lost zone, ZapRAID retrieves all available zones in the same segment from other available drives into memory, and reconstructs the stripe groups that cover the lost zone. It examines the block metadata of the available chunks and identifies the chunks from the same stripe. It then decodes the lost chunks for each stripe independently. After repairing all the stripes in a segment, ZapRAID writes the recovered zone to the new drive.
3.6 Discussion
We now discuss some open issues of our ZapRAID design.
Choice of the write request size threshold in hybrid data management. Recall that ZapRAID compares the size of a write request against a threshold to determine if the write request is assigned to a small-chunk or large-chunk open segment (§3.3). Currently, the threshold is set to $C_{l}$. The choice of the threshold poses different performance trade-offs. A larger threshold implies that more writes will go to small-chunk segments. Full small-chunk stripes are formed faster, and hence small writes are also acknowledged faster (i.e., lower response time). However, the write throughput is lower since writing small-chunk segments is slower than writing large-chunk segments (based on the results in Figure 2 for different write request sizes). For simplicity, we assume that the threshold is equal to $C_{l}$ (i.e., if a write request size is smaller than $C_{l}$, it is assigned to a small-chunk segment; otherwise, it is assigned to a large-chunk segment).
Relationships between the ZapRAID design and the ZNS SSD model. While our motivating experiments and evaluation are based on the ZN540 model [70], our ZapRAID design does not heavily depend on a specific ZNS SSD model. For example, some of our key design features, such as the group-based data layout and crash recovery management, apply to all ZNS SSD models.
ZapRAID is mainly based on two design principles: (i) Zone Append outperforms Zone Write in one or a few open zones for small writes, and (ii) a combination of Zone Append and Zone Write can be used for hybrid data management when there are multiple open zones. We note that ZapRAID now only uses Zone Append for a single open segment in hybrid data management since the improvement of Zone Append is limited for a large number of open zones, and we conjecture that the reason is due to the computational overhead of the firmware. Nevertheless, if the computational overhead can be mitigated in future ZNS SSD models, ZapRAID can issue Zone Append to more open segments to better exploit intra-zone parallelism; this can be readily achieved with parameter reconfigurations based on our current design.
Our ZapRAID design is also applicable to different zone capacities of ZNS SSDs. For smaller zone capacities, the data and footer regions become smaller, yet the group-based data layout remains applicable. For example, when the zone capacity is 96 MiB [49] (or equivalently, 24,576 blocks), for 4 KiB chunks, the header, data, and footer regions occupy 1 block, 24,455 blocks, and 120 blocks, respectively (based on the calculation of §3.1). If the stripe group size is 256, there are still many stripe groups (i.e., 96 groups) for intra-zone parallelism.
Note that the ZN540 model has been widely studied by a large body of prior work in the literature (e.g., [4, 31, 35, 60, 16]) and is a representative platform for current ZNS SSD research.
4 Implementation
Since ZNS SSDs offer low latency and high throughput, the overhead of the Linux kernel I/O stack (e.g., high latencies of interrupts, locks, and system calls) is no longer negligible [36]. To reduce I/O latency, we implement ZapRAID using the Storage Performance Development Kit (SPDK) [63]. SPDK supports polling-based and lockless I/Os. It also avoids system calls by moving I/O drivers to the user space.
Specifically, we prototype ZapRAID as a user-space block device module in C++ with around 9.7 K LoC based on SPDK. ZapRAID exports a block interface to be compatible with general applications. Our prototype decomposes request handling into seven handlers, namely dispatch, device I/O, completion, indexing, encoding, segment state tracking, and cleaning.
Dispatch handler. The dispatch handler is responsible for processing write requests and updating the header and footer regions. For write requests, the dispatch handler assigns the written blocks to segments and submits the I/O events for the written blocks to the device I/O handler. If a block fills a stripe in a segment, the dispatch handler also sends a parity generation event to the encoding handler, which computes parity chunks and submits the I/O events for the parity chunks to the device I/O handler. Recall that the writes of the first stripe of a stripe group need to wait for the previous stripe group to be completely persisted (§3.5); in this case, the dispatch handler will suspend the request and retry later until the segment state is updated by the segment state tracking handler (see below).
For read requests, the dispatch handler first sends an L2P table query event to the indexing handler, which asynchronously notifies the dispatch handler to issue the read requests upon retrieving PBAs for the requested blocks. The dispatch handler then submits the I/O events to the device I/O handler to retrieve the requested blocks from storage.
Device I/O handler. The device I/O handler is responsible for submitting block I/O requests to the drives. It receives I/O events from other handlers and submits the corresponding I/O requests to the drives. It is also responsible for checking the completion of each block I/O request from the drives and notifying the completion handler (see below).
Completion handler. The completion handler tracks the completion state of each I/O request. It receives the notification of completion of a block I/O event from the device I/O handler. If all the I/O events for a request are completed, the completion handler notifies the upper-layer application. The application can register a callback function for each request. For write requests, the completion handler first sends an L2P table update event to the indexing handler, which triggers the callback function after updating the L2P table. For read requests, the completion handler directly triggers the callback function.
If a block is lost or unavailable during a read request, the completion handler issues a degraded read to the block. It examines the compact stripe table to identify the PBAs of the available blocks of the same stripe. It then sends the I/O events to the device I/O handler to retrieve the available blocks. After all available blocks are retrieved, the completion handler decodes the lost block and marks the read request as completed.
Indexing handler. The indexing handler is responsible for the L2P table maintenance. It receives L2P table query events from the dispatch handler and returns the PBAs of requested blocks for read requests. It also receives L2P table update events from the completion handler and the cleaning handler. If an L2P table update event is from the completion handler, it acknowledges the write requests to the application on behalf of the completion handler after the L2P table is updated. In addition, the indexing handler is responsible for the offloading of L2P table entries if the L2P table size reaches the memory size limit (§3.1).
Encoding handler. The encoding handler is responsible for parity generation. When it receives a notification from the dispatch handler that a stripe has $k$ data chunks, it generates parity chunks. It then submits I/O events for the parity chunks to the device I/O handler. We handle parity chunk generation outside of the normal I/O requests to mitigate interference on normal I/O requests.
[…]
In this work, we do not further consider sophisticated garbage collection algorithms, which need to address the key design issues of (i) when to trigger garbage collection, (ii) which segment to clean for garbage collection, and (iii) where to rewrite the valid data. We argue that such design issues are orthogonal to our current focus on leveraging Zone Append and Zone Write for a high-performance RAID design. We pose the design of new garbage collection algorithms as future work.
Thread assignment. We assign the seven handlers to SPDK threads, which exchange events via spdk_thread_send_msg and event polling. Currently, we co-locate the dispatch handler, completion handler, segment state tracking handler, and cleaning handler in the main SPDK thread. We also assign the device I/O handler, indexing handler, and encoding handler separately to three different SPDK threads for parallelism, such that our prototype can saturate the write bandwidth of our testbed.
5 Evaluation
We evaluate ZapRAID through real-application experiments, trace-driven experiments, and microbenchmarks.
5.1 Methodology
We use a server that runs Ubuntu 22.04 LTS with Linux kernel 5.15. It has a 16-core Intel Xeon Silver 4215 2.5 GHz CPU and 96 GiB DRAM. It is attached with four 4 TiB Western Digital Ultrastar DC ZN540 ZNS SSDs [70]. Each SSD has 3,690 zones, with a zone capacity of 1,077 MiB each. We format each SSD with a logical block size of 4 KiB and a block metadata size (in the out-of-band area) of 64 bytes. By default, we focus on (3+1)-RAID-5, yet we also consider other RAID schemes and a larger-scale emulated ZNS SSD array based on FEMU [39]5.4).
We mainly consider two baselines: ZoneWrite-Only and ZoneAppend-Only. ZoneWrite-Only always uses Zone Write for all writes (i.e., $G=1$), while ZoneAppend-Only always uses Zone Append for all writes and does not employ a group-based data layout (i.e., $G=S$) (§3.2). For fair comparisons, when there are multiple open segments, ZoneWrite-Only and ZoneAppend-Only also write small chunks to small-chunk segments and large chunks to large-chunk segments, only using Zone Write and Zone Append, respectively.
We also compare ZapRAID with a simplified version of RAIZN [31], which we implement as a user-space block device based on SPDK (referred to as RAIZN-SPDK). Note that the open-source version of RAIZN [57] is currently implemented as a kernel device mapper and exports a ZNS interface as a logical ZNS SSD volume, so we re-implement RAIZN as a user-space block device for fair comparisons with ZapRAID. Specifically, the original RAIZN reserves at least three additional metadata zones (in addition to regular zones for data storage) in each drive, including one zone for appending partial parity updates before acknowledging each write request, one zone for appending metadata to ensure write atomicity, and at least one zone for the garbage collection of metadata zones. It caches the written data in an in-memory stripe buffer and writes the partial parity updates to the partial-parity-dedicated metadata zone, so as to support crash recovery even if a system crash happens before the whole stripe is written. When the stripe buffer is full, RAIZN generates and writes the complete parity chunks to the regular zones, and discards the corresponding partial parity updates. RAIZN differs from ZoneWrite-Only, ZoneAppend-Only, and ZapRAID in that it acknowledges each write request after the corresponding partial parity updates are appended to the dedicated metadata zone, while the other three schemes leverage concurrent write requests to form full stripes before acknowledging the write requests.
In our RAIZN-SPDK implementation, we focus on achieving high performance. We simplify our implementation by removing the metadata zones reserved for write atomicity and garbage collection. RAIZN-SPDK also reserves two metadata zones for high performance. Specifically, it writes partial parity updates to one of the two metadata zones. When the metadata zone becomes full, RAIZN-SPDK resets the zone and writes the partial parity updates to another metadata zone in parallel. For high performance, RAIZN-SPDK uses Zone Append to write partial parity updates to a metadata zone as described in [31]. When there are multiple open segments, RAIZN-SPDK uses Zone Write to write small chunks to small-chunk segments and large chunks to large-chunk segments, as in ZoneWrite-Only, while using Zone Append to write partial parity updates of all segments to a metadata zone.
We also attempt to fairly compare ZapRAID with the conventional RAID array that runs on traditional block-interface SSDs. However, the firmware on our ZN540 SSDs is locked and cannot be directly modified, so we cannot make comparisons on the same hardware as in [31]. Thus, we leverage FEMU-emulated SSDs [39] for comparisons and build the conventional RAID array with mdadm [47] (see Exp#10 for details).
Our experiments consider the configurations with a single open segment and multiple open segments. If there is a single open segment, ZapRAID always issues writes via Zone Append and applies the group-based data layout; if there are multiple open segments, ZapRAID enables hybrid segment management and issues writes via a combination of Zone Append and Zone Write (the exact configurations are further explained in individual experiments).
For each experiment, we report the average results over five runs and include the 95% confidence intervals based on Student’s t-distribution (note that some confidence intervals may be invisible due to negligible deviations across runs).
Our evaluation is driven by the following research questions:
  • How does ZapRAID perform on real applications in terms of writes, degraded reads, and memory usage for index management, when compared to ZoneWrite-Only and ZoneAppend-Only? (RQ1)
  • How does ZapRAID perform under realistic cloud storage traces that are composed of a mix of small and large writes? (RQ2)
  • How does ZapRAID perform across different aspects under microbenchmarks, including writes, normal reads, degraded reads, crash recovery, and full-drive recovery? (RQ3)
5.2 Real-Application Experiments (RQ1)
We start our evaluation with real applications to demonstrate the end-to-end performance benefits of ZapRAID, in comparison with ZoneWrite-Only and ZoneAppend-Only (similar benchmarks are also reported in [31]). We exclude RAIZN-SPDK, as it shows poor performance from our microbenchmarks (§5.4).
To host Linux applications, we export our ZapRAID prototype as an NVMe-over-Fabrics (NVMe-oF) target with TCP transport [64] and deploy the NVMe-oF host in the same testbed machine, so that the prototype serves as a Linux block device. We allocate the block device with a logical space capacity of 1 TiB, format the block device as an EXT4 file system, and run different types of applications atop the file system. Since the application write traffic may exceed the logical capacity, we reserve an additional 25% of the physical capacity for garbage collection.
We evaluate ZapRAID’s performance on both single-segment and multi-segment settings. For multi-segment settings, we vary the numbers of small-chunk and large-chunk segments (§3.3). We fix the small- and large-chunk sizes at $C_{s}=8$ KiB and $C_{l}=16$ KiB, respectively, and keep the total number of open segments as four (i.e., $N_{s}+N_{l}=4$). Here, we focus on four representative settings of open segments: (i) a single open segment with a chunk size of 4 KiB, (ii) a single open segment with a chunk size of 16 KiB, (iii) four open segments with $(N_{s},N_{l})=(2,2)$ (i.e., two small-chunk segments and two large-chunk segments), and (iv) four open segments with $(N_{s},N_{l})=(1,3)$ (i.e., one small-chunk segment and three large-chunk segments). We also present extensive evaluation results of various open-segment configurations in microbenchmarks in §5.4.
Workloads. We evaluate three types of application workloads: file writes, key-value workloads, and concurrent workloads.
File writes. We evaluate ZapRAID in file writes using Sysbench (v1.0.20) [66] to generate continuous write patterns. We first create 128 files of 800 MiB each (totaling 100 GiB). We then use Sysbench to issue continuous write requests with request sizes of 4 KiB, 8 KiB, 16 KiB, 32 KiB, and 64 KiB. For each request size, we run the workload for 10 minutes. Unless otherwise specified, we use one thread to issue writes in asynchronous mode with direct I/O, and each run ends with an fsync to ensure data persistence. Before each run, we clear and reformat the block device as an EXT4 file system. This workload stresses small, frequent writes and highlights both the throughput and metadata efficiency of different schemes.
Key-value store workloads. We evaluate ZapRAID when it hosts RocksDB (v9.7.2) [58] and use its benchmarking tool db_bench [53] to issue key-value operations. We focus on two settings of the fillrandom workload: (i) the default fillrandom, which inserts randomly ordered keys into an empty database and triggers background compaction, and (ii) fillrandom with compaction disabled (bulk loading), which sequentially writes data without compaction overhead. We configure the database with approximately 1.64 million key-value pairs, each with a value size of 64 KiB, and disable compression. We use eight threads per client and run a total of eight clients concurrently, resulting in 64 writer threads in total. The workload writes about 100 GiB of data in total, evenly distributed across all clients. For each run, we reset the volume and then issue one of the two fillrandom workloads. These two workload settings capture both the compaction overhead of LSM-tree maintenance and the pure bulk-loading behavior without compaction.
Concurrent workloads. We combine two types of write-intensive workloads with different write patterns to evaluate ZapRAID. Specifically, we use Sysbench to generate small file writes, while simultaneously hosting MySQL and running the oltp_write_only benchmark, which issues large writes with fewer than 10% reads. This setting emulates a virtualized environment in which multiple applications coexist in an SSD RAID array [14]. For the small file writes, we use Sysbench to continuously issue 4 KiB write requests. For the oltp_write_only workload, we use Sysbench to populate the MySQL database with 8 tables and 10 million rows, and then execute DELETE, INSERT, and UPDATE queries using 64 threads. Before each run, we clear and reformat the block device as an EXT4 file system, and then execute both benchmarks simultaneously for 10 minutes.
Refer to caption Refer to caption Refer to caption Refer to caption
(a) File writes, single (4 KiB) (b) File writes, single (16 KiB) (c) File writes, multiple $(N_{s},N_{l})=(2,2)$ (d) File writes, multiple $(N_{s},N_{l})=(1,3)$
Refer to caption Refer to caption Refer to caption Refer to caption
(e) Key-value (with compaction), single (f) Key-value (with compaction), multiple (g) Key-value (bulk load, no compaction), single (h) Key-value (bulk load, no compaction), multiple
Refer to caption Refer to caption Refer to caption Refer to caption
(i) Concurrent: file writes (4 KiB), single (j) Concurrent: file writes (4 KiB), multiple (k) Concurrent: oltp_ write_only, single (l) Concurrent: oltp_ write_only, multiple
Figure 6: Exp#1 (Write performance): file writes (figures (a)-(d)), key-value store workloads (figures (e)-(h)), and concurrent workloads (figures (i)-(l)).
Exp#1 (Write performance). We first evaluate the write performance of ZapRAID and compare it against ZoneWrite-Only and ZoneAppend-Only under different open-segment settings.
File writes. Figures 6(a)-6(d) present the throughput results of sequential file writes under different open-segment settings. For a single open segment with 4 KiB chunks (Figure 6(a)), ZapRAID achieves throughput comparable to ZoneAppend, while outperforming ZoneWrite-Only by 38.6%. At larger request sizes, the performance gap between ZapRAID and ZoneWrite-Only increases significantly, reaching 60.9%, 73.0%, 71.2%, and 76.8% at 8 KiB, 16 KiB, 32 KiB, and 64 KiB, respectively. The gains mainly stem from ZapRAID’s ability to exploit intra-zone parallelism under a small chunk size, which is consistent with our observation in Figure 2(a) in §2.2 for a single open segment with 4 KiB chunks.
At small request sizes (e.g., 4 KiB), ZapRAID’s improvement is relatively modest due to the system overhead from file system metadata updates, as well as kernel I/O stack and NVMe-oF’s TCP stack operations. As the request size becomes larger, such overhead becomes amortized. Since the chunk size is 4 KiB, large requests are still issued to small chunks. ZapRAID can deliver progressively larger throughput gains over ZoneWrite-Only via intra-zone parallelism.
Settings Request sizes Findings Reasons
Single segment, 4 KiB chunks 4-8 KiB Modest improvements over ZoneWrite-Only Reasons A and B
16-64 KiB Significant improvements over ZoneWrite-Only Reason A
Single segment, 16 KiB chunks 4-8 KiB Marginal difference Reason C
16-64 KiB
Multiple segments, $(N_{s},N_{l})=(2,2)$ 4-8 KiB Marginal difference Reason D
16-64 KiB Marginal difference Reason C
Multiple segments, $(N_{s},N_{l})=(1,3)$ 4-8 KiB Modest improvements over ZoneWrite-Only Reasons A and B
16-64 KiB Marginal difference Reason C
Reason A: Exploiting intra-zone parallelism with small chunks; see Figures 2(a) and 2(b) in §2.2.
Reason B: System overhead of file system metadata updates, I/O stacks, and NVMe-oF’s TCP stacks.
Reason C: Large writes have similar throughput in Zone Append and Zone Write; see Figure 2(c) in §2.2.
Reason D: Small writes are distributed across zones and have similar throughput in Zone Append and Zone Write; see Figures 2(a) and 2(b) in §2.2.
Table 1: Exp#1 (Write performance). Explanations of the throughput results for file writes across different request sizes and segment settings.
For a single open segment with 16 KiB chunks and multiple open segments with $(N_{s},N_{l})=(2,2)$ (Figures 6(b) and 6(c), respectively), the performance differences among the three schemes are small. For the former, the 4 KiB writes are absorbed into 16 KiB chunks and issued as large writes, which are handled comparably by Zone Append and Zone Write (Figure 2(c) in §2.2); for the latter, the 4 KiB writes are distributed across two zones, where Zone Append and Zone Write deliver similar performance (Figure 2(a) in §2.2). For multiple open segments with $(N_{s},N_{l})=(1,3)$ (Figure 6(d)), ZapRAID shows a modest advantage at small request sizes (subject to the system overhead), outperforming ZoneWrite-Only by 20.2% and 16.6% at 4 KiB and 8 KiB, respectively, while the difference narrows at large request sizes since large writes yield comparable throughput across schemes.
Table 1 summarizes the reasons above to provide a consolidated view across different request sizes and open segment settings.
Note that across all settings (Figures 6(a)-6(d)), small requests of 4 KiB and 8 KiB can only achieve the throughput of at most 1,061.50 MiB/s and 1,789.36 MiB/s, respectively. The throughput is constrained by the overhead of file system metadata updates, kernel I/O stack operations, and TCP stack operations in NVMe-oF. In Exp#1, each request must traverse both the file system and TCP/IP stacks, which incur substantial processing overhead. In contrast, Exp#5 (§5.4) issues requests directly to the SPDK target without traversing the TCP stack, and the throughput can reach up to 1,599 MiB/s and 3,012 MiB/s under 4 KiB and 8 KiB writes, respectively. The TCP overhead in NVMe-oF, even with TCP loopback, has also been reported by the SPDK community [67]. Such overhead becomes less significant for large request sizes.
Key-value store workloads. Figures 6(e)-6(h) show the throughput results of RocksDB using db_bench under two configurations: the fillrandom workload with compaction enabled and the bulk-loading workload without compaction.
Under fillrandom with compaction (Figures 6(e) and 6(f)), ZapRAID outperforms ZoneWrite-Only by 41.7% under a single open segment with 4 KiB chunks, while the three schemes perform nearly the same as under a single open segment with 16 KiB chunks (Figure 6(e)). For multiple open segments with $(N_{s},N_{l})=(2,2)$ and $(N_{s},N_{l})=(1,3)$, the throughput difference across the three schemes remains small (Figure 6(f)).
For bulk loading without compaction (Figures 6(g) and 6(h)), ZapRAID shows larger benefits at small request sizes. It achieves 51.7% higher throughput than ZoneWrite-Only under a single open segment with 4 KiB chunks, while the advantage narrows to 5.7% under a single open segment with 16 KiB chunks. As with the compaction-enabled workload, the multiple open segments with $(N_{s},N_{l})=(2,2)$ and $(N_{s},N_{l})=(1,3)$ yield comparable throughput across all schemes. Since compaction is disabled during bulk loading, read overheads are eliminated, making the performance improvement of ZapRAID more pronounced.
We observe that the performance gaps across the three schemes are small for the configurations with a large chunk size (e.g., 16 KiB) or sufficient open segments (e.g., four). RocksDB exhibits both small writes (e.g., write-ahead log (WAL) writes) and large writes (e.g., flushes and compactions), while large writes dominate. Both Zone Write and Zone Append primitives are expected to have close performance in such configurations (see Figure 2 in §2.2).
Concurrent workloads. Figures 6(i) and 6(j) show the write throughput results for small file writes in concurrent workloads. ZapRAID achieves 26.3% and 34.5% higher throughput than ZoneWrite-Only under a single open segment with 4 KiB chunks (Figure 6(i)) and four open segments with $(N_{s},N_{l})=(1,3)$ (Figure 6(j)), respectively. In both settings, all small writes are directed to the single small-chunk segment. ZapRAID outperforms ZoneWrite-Only by issuing Zone Append. On the other hand, for a single open segment with 16 KiB chunks and four open segments with $(N_{s},N_{l})=(2,2)$, all three schemes have similar performance, due to the same reasons as explained in the file writes workloads (see above).
Figures 6(k) and 6(l) depict the throughput of the oltp_write_only benchmark in transactions per second (TPS). For a single open segment with 4 KiB chunks (Figure 6(k)), ZapRAID achieves 32.9% higher throughput than ZoneWrite-Only and has similar throughput to ZoneAppend-Only. For other open-segment settings, all three schemes show similar performance. In oltp_write_only, where large writes dominate, we expect similar performance across all three schemes for the configurations with a large chunk size and sufficient open segments (see the explanations above for key-value store workloads).
Exp#2 (Degraded reads). We evaluate degraded read performance by first loading data into the system, failing one disk to emulate degraded mode, and finally issuing read requests based on the file writes and key-value store workloads. We consider two workloads, derived from the file writes and key-value store workloads, respectively: (i) Sysbench with readrandom and (ii) db_bench with readrandom. For the former, we load a 100 GiB dataset using Sysbench, fail one device, and perform degraded random reads continuously for 10 minutes using Sysbench in readrandom mode. For the latter, we use db_bench to first load a 100 GiB database with 64 KiB values, fail one device, and then issue 20,000 degraded readrandom requests using db_bench.
[…]
Refer to caption Refer to caption Refer to caption Refer to caption
(a) Sysbench, single (b) Sysbench, multiple (c) db_bench, single (d) db_bench, multiple
Figure 7: Exp#2 (Degraded reads): Sysbench (figures (a) and (b)) and db_bench (figures (c) and (d)).
[…]
File writes Single-4 KiB Single-16 KiB $(N_{s},N_{l})=(2,2)$ $(N_{s},N_{l})=(1,3)$
(ZoneAppend-Only / ZapRAID)
L2P Table 100.00 / 100.00 100.00 / 100.00 100.00 / 100.00 100.00 / 100.00
Segment table 0.0008 / 0.0008 0.0008 / 0.0008 0.0008 / 0.0008 0.0008 / 0.0008
Compact stripe table 106.68 / 35.56 26.67 / 8.89 26.67 / 0.00 26.67 / 0.00
Total 206.68 / 135.56 126.67 / 108.89 126.67 / 100.00 126.67 / 100.00
Key-value store Single-4 KiB Single-16 KiB $(N_{s},N_{l})=(2,2)$ $(N_{s},N_{l})=(1,3)$
(ZoneAppend-Only / ZapRAID)
L2P table 77.55 / 77.55 77.55 / 77.55 77.55 / 77.55 77.55 / 77.55
Segment table 0.0006 / 0.0006 0.0006 / 0.0006 0.0006 / 0.0006 0.0006 / 0.0006
Compact stripe table 81.58 / 27.19 20.40 / 6.80 20.40 / 0.00 20.40 / 0.00
Total 159.13 / 104.74 97.95 / 84.35 97.95 / 77.55 97.95 / 77.55
Table 2: Exp#3 (Memory usage for index management). Breakdown of memory usage (MiB) for index management in ZoneAppend-Only and ZapRAID under the file writes and key-value store workloads.
Exp#3 (Memory usage for index management). To evaluate memory usage for index management, we consider the file writes and key-value store workloads. We assume that the L2P table is entirely in memory, without being offloaded (§3.1). Table 2 reports the breakdown of memory usage for index management (including the L2P table, segment table, and compact stripe table) and the total memory usage.
For file writes, we run Sysbench with a request size of 64 KiB to generate a 100 GiB dataset, while for key-value store workloads, we run db_bench with compaction enabled to write a total of 100 GiB of data across eight clients with 64 KiB requests. We compare ZoneAppend-Only and ZapRAID, since they both maintain stripe-to-chunk mappings under Zone Append, while ZoneWrite-Only does not issue Zone Append and hence has no compact stripe table memory consumption. The workload settings are identical to those used in Exp#1.
For file writes, the L2P table and segment table sizes stay constant across configurations (100.00 MiB and 0.0008 MiB, respectively), as they depend only on the total volume of data written. The differences come from the compact stripe table. For a single open segment with 4 KiB chunks, the compact stripe table decreases from 106.68 MiB (ZoneAppend-Only) to 35.56 MiB (ZapRAID) (i.e., 66.7% reduction), so ZapRAID reduces the total memory usage of ZoneAppend-Only from 206.68 MiB to 135.56 MiB (34.4%). We emphasize that ZapRAID’s memory reduction (66.7%) of the compact stripe table compared to ZoneAppend-Only is critical, since it significantly reduces the degraded read overhead (Exp#2).
ZapRAID also shows memory usage reduction in other settings. For a single open segment with 16 KiB chunks, the compact stripe table drops from 26.67 MiB to 8.89 MiB (66.7%), and ZapRAID reduces the total memory usage of ZoneAppend-Only from 126.67 MiB to 108.89 MiB (14.0%). For multiple open segments with $(N_{s},N_{l})=(2,2)$ and $(N_{s},N_{l})=(1,3)$, ZapRAID directs all 64 KiB writes to large segments, so it reverts to ZoneWrite and does not require a compact stripe table. Thus, ZapRAID reduces the compact stripe table size from 26.67 MiB to 0.00 MiB, and the total memory usage is reduced from 126.67 MiB to 100.00 MiB (21.1%), compared to ZoneAppend-Only.
For key-value store workloads, the trends are similar. The L2P and segment table sizes remain fixed at 77.55 MiB and 0.0006 MiB, respectively, while the compact stripe table and total memory usage show clear savings in ZapRAID. For example, for a single open segment with 4 KiB chunks, ZapRAID reduces the compact stripe table size from 81.58 MiB to 27.19 MiB (66.7%), reducing the total memory usage from 159.13 MiB to 104.74 MiB (34.2%), compared to ZoneAppend-Only.
Summary of real-application experiments. Compared to ZoneWrite-Only, ZapRAID excels in small chunk sizes and increases the write throughput by up to 76.8% (in 64 KiB writes under Sysbench). A trade-off is that ZapRAID introduces a compact stripe table, but it limits the memory usage of the compact stripe table (as well as the degraded read overhead) through the group-based data layout.
Compared to ZoneAppend-Only, ZapRAID achieves nearly identical write throughput (Exp#1), while significantly improving degraded read performance (Exp#2) and reducing memory usage (Exp#3). For example, in Sysbench file writes under a single open segment with 4 KiB chunks, ZapRAID’s write throughput is within 2.42% of ZoneAppend-Only’s, while reducing the compact stripe table size by 66.7% and the total memory usage for index management by 34.4%. Similarly, in the db_bench fillrandom workload under a single segment with 4 KiB chunks, ZapRAID’s write throughput is within 2.20% of ZoneAppend-Only’s, while reducing the compact stripe table size by 66.7% and the total memory usage for index management by 34.2%. By limiting the compact stripe table size, ZapRAID also significantly outperforms ZoneAppend-Only in degraded reads.
5.3 Trace-Driven Experiments (RQ2)
We evaluate via trace-driven evaluation ZapRAID on real-world cloud block storage workloads using the block I/O traces from Alibaba Cloud [41]. We configure the logical space capacity as 500 GiB. Our evaluation considers four settings of open segments: (i) a single open segment with a chunk size of 4 KiB, (ii) a single open segment with a chunk size of 16 KiB, (iii) four open segments with $(N_{s},N_{l})=(2,2)$, and (iv) four open segments with $(N_{s},N_{l})=(1,3)$; for (iii) and (iv), we set $C_{s}=$ 8 KiB and $C_{l}=$ 16 KiB.
Exp#4 (Cloud block storage workloads). The original Alibaba Cloud traces cover 1,000 volumes [41]. We identify that 684 out of the 1,000 volumes have more than 60% of small write requests of no more than 4 KiB and more than 25% of large write requests of at least 16 KiB (§2.2). Among the 684 volumes, we choose 20 of them whose amounts of write traffic range between 200 GiB and 240 GiB. Note that the selected 20 volumes are dominated by writes: in 18 out of 20 volumes, more than 88.3% of requests are writes, while in all volumes, more than 67.0% of requests are writes. We conduct our evaluation over five runs; in each run, we evaluate the average throughput and 95th-percentile latency of all read and write operations when running each scheme over the 20 selected volumes.
Figures 8(a) and 8(b) show the throughput results. For a single open segment with 4 KiB and 16 KiB chunks (Figure 8(a)), the throughput of ZapRAID is 69.4% and 6.4% higher than that of ZoneWrite-Only, respectively, and is close to that of ZoneAppend-Only; For multiple open segments $(N_{s},N_{l})=(2,2)$, all three schemes have similar throughput, while for $(N_{s},N_{l})=(1,3)$, ZapRAID has 25.3% higher throughput than ZoneWrite-Only and similar throughput to ZoneAppend-Only.
Refer to caption Refer to caption Refer to caption Refer to caption
(a) Throughput (single segment) (b) Throughput (multiple segments) (c) p95 latency (single segment) (d) p95 latency (multiple segments)
Figure 8: Exp#4 (Cloud block storage workloads). Average throughput and 95 latency results of all 20 volumes in Alibaba Cloud traces.
Figures 8(c) and 8(d) show the 95th-percentile latencies. For a single open segment with 4 KiB and 16 KiB chunks (Figure 8(c)), ZapRAID has 56.6% lower and 39.6% higher 95th-percentile latencies than ZoneWrite-Only under the chunk sizes of 4 KiB and 16 KiB, respectively, and comparable 95th-percentile latencies with ZoneAppend-Only. For multiple open segments (Figure 8(d)), all three schemes have similar latencies for $(N_{s},N_{l})=(2,2)$, while ZoneWrite-Only has a higher 95th-percentile latency for $(N_{s},N_{l})=(1,3)$ as its performance is bottlenecked by the small writes that are forwarded to the only small-chunk open segment.
Volume No. 1 2 3 4 5 6 7 8 9 10
Ratios (%) Small writes 83.35 83.06 83.30 83.56 83.29 83.50 83.03 83.36 82.66 82.98
Large writes 1.63 1.64 1.67 1.70 1.75 1.76 1.77 1.78 1.83 1.86
Thpt (MiB/s) (4 KiB chunks) ZoneWrite-Only 1,023.7 947.0 1,035.1 1,001.5 937.5 981.5 1,035.3 968.2 986.2 988.7
ZapRAID 1,740.6 1,711.3 1,702.4 1,699.1 1,660.1 1,657.9 1,612.7 1,619.4 1,681.0 1,636.3
Thpt (MiB/s) $($$N_{s},N_{l}$)=(1,3)) ZoneWrite-Only 2,486.4 2,341.2 2,323.1 2,401.3 2,403.7 2,333.7 2,311.0 2,360.3 2,344.8 2,410.5
ZapRAID 3,175.9 3,225.1 3,271.9 3,193.7 3,183.1 3,227.5 3,231.6 3,205.8 3,249.7 3,298.2
Volume No. 11 12 13 14 15 16 17 18 19 20
Ratios (%) Small writes 82.96 86.68 84.00 81.58 80.98 81.41 72.45 73.85 65.78 63.05
Large writes 1.92 1.95 3.42 3.70 4.49 10.33 16.83 20.95 21.11 24.89
Thpt (MiB/s) (4 KiB chunks) ZoneWrite-Only 948.4 1,007.2 982.2 995.6 989.2 972.1 970.4 993.0 1,004.0 971.5
ZapRAID 1,627.3 1,710.9 1,634.0 1,645.8 1,699.1 1,666.3 1,684.8 1,736.4 1,656.4 1,665.6
Thpt (MiB/s) $($$N_{s},N_{l}$)=(1,3)) ZoneWrite-Only 2,372.7 2,399.1 2,412.2 2,439.7 2,593.6 2,637.1 2,703.2 2,773.4 2,862.1 2,825.4
ZapRAID 3,175.9 3,225.1 3,271.9 3,193.7 3,183.1 3,227.5 3,231.6 3,205.8 3,249.7 3,298.2
Table 3: Exp#4 (Cloud block storage workloads). Statistics of the 20 individual volumes in Alibaba Cloud traces.
We also show the statistics of individual volumes by presenting the ratios of numbers of small writes (i.e., 4 KiB) and large writes (i.e., at least 16 KiB) over the total number of write requests as well as the throughput of each volume. For brevity, since all three schemes have close throughput under one single open segment with 16 KiB chunks and four open segments with $(N_{s},N_{l})=(2,2)$, we only present the results of one single open segment with 4 KiB chunks and four open segments with $(N_{s},N_{l})=(1,3)$. Also, we only focus on the results under ZoneWrite-Only and ZapRAID to show their variations across the volumes.
Table 3 shows the results. We sort the volumes by the ratio of large write requests. For one single open segments under 4 KiB chunks, the throughput is close over all volumes for each scheme. For four open segments with $(N_{s},N_{l})=(1,3)$, ZoneWrite-Only shows lower throughput for a lower ratio of large writes; for example, the throughput results of Volumes 2 and 20 are 2,341.2 MiB/s and 2,825.4 MiB/s with large-write ratios of 1.64% and 24.89%, respectively. The reason is that with fewer large writes, more write requests are written to the only $N_{s}=1$ small-chunk open segment, and ZoneWrite-Only fails to utilize intra-zone parallelism to issue writes efficiently. In contrast, ZapRAID maintains higher throughput with an increase of 14.7-40.8% compared with ZoneWrite-Only under $(N_{s},N_{l})=(1,3)$, across all 20 volumes.
5.4 Microbenchmarks (RQ3)
We examine the performance of ZapRAID in writes, normal reads, degraded reads, crash recovery, and full-drive recovery. In our default setup, we configure the logical space capacity as 200 GiB and set $G=256$ for ZapRAID in its group-based data layout. First, we focus on one open segment with various chunk sizes in Exp#5-Exp#10. We do not consider garbage collection (by limiting the total write size to be less than the logical space capacity) and keep the L2P table entirely in memory. We also consider the impact of various group sizes (Exp#7), the impact of RAID schemes (Exp#8), as well as the crash recovery and full-drive recovery performance with different logical space capacities (Exp#9). Furthermore, we consider multiple open segments in Exp#11-Exp#13, including the write performance under multiple open segments (Exp#11), the overhead of garbage collection (Exp#12), and the overhead of L2P table offloading (Exp#13).
Exp#5 (Write performance). We first evaluate the write performance of ZapRAID on a single open segment. We use Flexible IO Tester (FIO) (v3.35) [17] to generate a write-only workload that issues random writes of 64 GiB of data. We set the queue depth as 64 to saturate the system parallelism (Exp#10). We compare ZapRAID with RAIZN-SPDK, ZoneWrite-Only, and ZoneAppend-Only.
First, we issue write requests of 4 KiB, 8 KiB, and 16 KiB, where the request size is the same as the chunk size. Figures 9(a)-9(c) show the throughput as well as the median and 95th-percentile latencies. Compared with ZoneWrite-Only, for 4 KiB and 8 KiB writes, ZapRAID achieves 72.8% and 77.2% higher write throughput, 43.8% and 44.5% lower median latency, and 36.4% and 34.1% lower 95th-percentile latency, respectively. This shows that ZapRAID improves the write performance by exploiting the intra-zone parallelism using Zone Append for small writes. On the other hand, for 16 KiB writes, ZapRAID has similar write throughput and median latency, but its 95th-percentile latency is $3.80\times$ compared with that of ZoneWrite-Only. A possible reason is that Zone Append has high computational overhead in the ZNS SSD firmware, which incurs higher delays in processing individual Zone Append commands.
[…]
Compared with ZoneAppend-Only, ZapRAID achieves similar throughput and latencies for all write sizes, even though ZapRAID needs to wait for the completion of all Zone Append commands in a stripe group before issuing the writes for the next stripe group. Note that ZapRAID has much less memory usage (§3.2) and much lower degraded read latency (Exp#7) than ZoneAppend-Only.
Compared with RAIZN-SPDK, all other three schemes show much higher performance on a single segment. The reasons are two-fold. First, RAIZN does not support Zone Append for intra-zone parallelism. Second, it uses partial parity updates to acknowledge individual write requests instead of issuing full-stripe writes as in the other three schemes. This limits the concurrency of write requests and hence the overall write performance. In Exp#11, we further compare ZapRAID with RAIZN-SPDK under multiple open segments.
Figure 9(d) shows the memory usage for stripe management per GiB of written data. Here, we focus on ZoneAppend-Only and ZapRAID, as they both need to maintain mappings between stripes and chunk locations under Zone Append (note that ZoneWrite-Only has no compact stripe table as it does not issue Zone Append). For 4 KiB writes, ZoneAppend-Only consumes 1,018.2 KiB of memory, while ZapRAID consumes 339.4 KiB only (i.e., 66.7% reduction) due to its compact stripe table design. For a (3+1)-RAID-5 array in which each drive has $Z=$ 3,690 zones and the zone capacity is 1,077 MiB, the maximum memory size for ZoneAppend-Only is $\frac{1018.2\times 3\times 3690\times 1077}{1024\times 1024\times 1024}=11.3$ GiB, while ZapRAID consumes 3.77 GiB only (§3.2). The result suggests that ZapRAID greatly reduces the memory footprints for stripe management compared with ZoneAppend-Only. Note that larger chunks imply a smaller compact stripe table since each segment has fewer chunks.
Refer to caption Refer to caption
(a) Write throughput (b) Median write latency
Refer to caption Refer to caption
(c) p95 write latency (d) Memory usage (per GiB of data) for compact stripe table
Figure 9: Exp#5 (Write performance on a single open segment). We show the results when the request size is equal to the chunk size. For figure (d), we show ZoneAppend-Only and ZapRAID only.
Refer to caption Refer to caption
(a) Write throughput (b) Median write latency
Refer to caption Refer to caption
(c) p95 write latency (d) Memory usage (per GiB of data) for compact stripe table
Figure 10: Exp#5 (Write performance on a single open segment). We show the results when the request size is fixed as 4 KiB.
We also issue write requests of a fixed size of 4 KiB, while varying the chunk size, to show the impact of the chunk size on write performance. Figure 10 shows the results. ZapRAID still maintains the high throughput and low median latencies for various chunk sizes, but its 95th-percentile latency for 16 KiB chunks is 50.7% higher than that for 8 KiB chunks. Similar results also appear in ZoneAppend-Only. The result indicates that Zone Append is mainly beneficial for small writes. It also justifies why we choose Zone Write over Zone Append in hybrid data management (§3.3).
Exp#6 (Normal and degraded read performance). We compare the performance of both normal reads and degraded reads in Log-RAID and ZapRAID. We vary the chunk size and set the read size to be the same as the chunk size. For each chunk size, we first fill up the 200 GiB logical address space to ensure that every logical block being read is indeed physically stored. We then use FIO to generate a read-only workload that issues random reads to 16 GiB of data using the randread option in FIO. We set the queue depth as one, so as to focus on the performance of individual read requests and exclude the interference among requests. We consider three cases: normal reads (NR), degraded reads under the static mapping in Log-RAID (DR-Log-RAID) (§2.3), and degraded reads under the group-based data layout in ZapRAID (DR-ZapRAID). Note that both Log-RAID and ZapRAID have the same workflow for normal reads. To evaluate degraded reads, we fail a drive and issue reads to the lost blocks of the failed drive.
Figure 11 shows the throughput as well as the median and 95th-percentile latencies for different read sizes (i.e., chunk sizes). All three read operations have less than 6% difference. The degraded reads of both Log-RAID and ZapRAID have slightly worse performance than the normal reads, since both systems retrieve multiple available chunks in parallel for decoding. Also, ZapRAID achieves comparable performance to Log-RAID, which uses static mapping.
Refer to caption Refer to caption Refer to caption
(a) Read throughput (b) Median read latency (c) p95 read latency
Figure 11: Exp#6 (Normal and degraded read performance). We consider normal reads (NR), degraded reads under static mapping in Log-RAID (DR-Log-RAID), and degraded reads under the group-based data layout in ZapRAID (DR-ZapRAID).
Exp#7 (Impact of stripe group size). We study the impact of the stripe group size $G$ on the write and read performance in ZapRAID. We vary $G$ from 4 to 4,096. For both writes and reads, we vary the chunk size and set the request size to be the same as the chunk size.
[…]
Refer to caption Refer to caption
(a) Write throughput for ZapRAID (b) Median degraded read latencies for different schemes
Figure 12: Exp#7 (Impact of stripe group size).
[…]
Exp#8 (Impact of RAID schemes). ZapRAID supports various RAID schemes. We configure RAID-0, RAID-01, RAID-4, RAID-5, and RAID-6 on the four ZNS SSDs (e.g., $k=2$ and $m=2$ in RAID-6). Here, we compare ZapRAID and ZoneWrite-Only. We fix the request size to be the same as the chunk size and vary the chunk size with a queue depth of 64.
Figure 13 depicts the results. ZapRAID increases the write throughput of ZoneWrite-Only by 71.5-72.1%, 75.8-76.4%, and 5.3-5.7% for 4 KiB, 8 KiB, and 16 KiB chunks, respectively. The difference in the write throughput among the RAID schemes is due to the different number of data chunks in each stripe (i.e., four, two, three, three, and two for RAID-0, RAID-01, RAID-4, RAID-5, and RAID-6, respectively). The results show that ZapRAID consistently increases the write throughput of ZoneWrite-Only under different RAID schemes.
Refer to caption Refer to caption Refer to caption
(a) 4 KiB chunks (b) 8 KiB chunks (c) 16 KiB chunks
Figure 13: Exp#8 (Impact of RAID schemes).
Exp#9 (Recovery performance). We evaluate the recovery performance of ZapRAID in two aspects: recovery from a system crash (§3.4) and recovery from a full-drive failure (§3.5). We sequentially write data to ZapRAID configured with a fixed size of logical space (varying from 100 GiB to 1,000 GiB), mimic each type of failure, and repeat the recovery procedure for five runs.
We first evaluate the crash recovery time of ZapRAID. We crash ZapRAID and rebuild it from the currently stored data in the drives. We report the average recovery time (i.e., the time spent on performing all recovery operations in §3.4). Figure 14(a) depicts the results for different chunk sizes. For the 100 GiB storage space, ZapRAID takes 2.38-2.60 s to recover the system state. The recovery time increases linearly with the logical space size. The reason is that for larger logical space, ZapRAID needs to recover the system state from more sealed segments by reading their footer regions. For example, for the 1,000 GiB storage space, the recovery time increases to 15.92-16.22 s. Nevertheless, the recovery overhead is limited. For example, for 4 KiB chunks, we only need to pay an additional (16.222.60)/(101)=1.51(16.22-2.60)/(10-1)=1.51 s for recovering the system state from each 100 GiB of additional logical space. Also, the crash recovery time remains almost the same for different chunk sizes, since most of the time is spent on reading the footer regions with a large read size, independent of the chunk size.
We next evaluate the full-drive recovery time of ZapRAID. We erase all data in one drive and recover the lost data in the same drive. We report the average recovery time as the total time spent on reading available data from existing drives, reconstructing the stripes, and writing the recovered data to the new drive. Figure 14(b) shows the results for different chunk sizes. The full-drive recovery time is proportional to the logical space size. For example, for 4 KiB chunks, ZapRAID takes 81.9 s and 813.2 s to recover the lost data for the 100 GiB and 1,000 GiB storage space, respectively. On average, for each 100 GiB storage space, ZapRAID spends $813.2/10=81.3$ s to recover the lost data. Note that the recovery time is smaller for larger chunk sizes since ZapRAID can write recovered data with a larger write size. For example, for 8 KiB and 16 KiB chunks, the recovery time is reduced by 18.1-22.2% and 22.1-23.8% compared with 4 KiB chunks, respectively.
Refer to caption Refer to caption
(a) Crash recovery (b) Full-drive recovery
Figure 14: Exp#9 (Recovery performance).
Exp#10 (Scalability). We examine the scalability of ZapRAID by evaluating its write performance under different values of queue depth and showing how it saturates the system parallelism with a high queue depth. We consider both real and FEMU-emulated [39] ZNS SSDs.
We first vary the queue depth in our testbed with the real ZNS SSDs. We use the write-only workload, vary the chunk size, and set the write size to be the same as the chunk size. We set the minimum queue depth as four to evaluate ZapRAID under concurrent writes.
[…]
We also examine the performance of the conventional RAID array based on FEMU-emulated block-interface SSDs. Similar to FEMU ZNS SSDs, we configure a (6+1)-RAID-5 array on seven FEMU black-box SSDs with Linux mdadm [47]. The configurations of flash blocks, channels, and chips are the same as FEMU ZNS SSDs, and each FEMU SSD has 16 GiB of physical space. We also keep the same page write latency and the same settings as the FEMU virtual machine. We export the RAID array as a block device with 64 GiB of logical space, so the overprovisioned space is about $(\frac{16\times 6}{64}-1)\times 100\%=50\%$. We use FIO to issue random writes of 64 GiB to the RAID array for various chunk sizes and queue depths. Figure 15(c) shows the results. Compared with ZapRAID in Figure 15(b), the throughput of the conventional RAID array is much lower (up to 200 MiB/s only). The results are expected, since updating a data chunk in the conventional RAID array issues a read-modify-write operation, which reads the current data chunk being modified and the current parity chunk, modifies the parity chunk, and writes back the new data chunk and new parity chunk. In contrast, ZapRAID issues concurrent write requests to form full stripes and avoids the overhead of parity updates.
Refer to caption Refer to caption Refer to caption
(a) ZapRAID on real ZNS SSDs (b) ZapRAID on FEMU ZNS SSDs (c) Conventional RAID on FEMU block-interface SSDs
Figure 15: Exp#10 (Scalability). The results in figures (a) and (b) are based on ZapRAID, and the results in figure (c) are based on the conventional RAID array.
Exp#11 (Multiple open segments). We evaluate the write performance ZapRAID on multiple open segments based on hybrid data management (§3.3) by varying the numbers of small-chunk and large-chunk segments (i.e., $N_{s}$ and $N_{l}$, respectively). As introduced in §5.2, we fix $C_{s}=$ 8 KiB, $C_{l}=$ 16 KiB, and the number of open segments as four. We consider different combinations of $N_{s}$ and $N_{l}$, such that $N_{s}+N_{l}=4$. Note that if $N_{s}=0$ (resp. $N_{l}=0$), all writes are issued to large-chunk (resp. small-chunk) segments.
We consider four workloads. The first three workloads issue 4 KiB, 8 KiB, and 16 KiB writes via FIO, respectively. Note that these workloads do not use all open segments since they write 4 KiB or 8 KiB requests to small-chunk segments only and 16 KiB requests to large-chunk segments only. The fourth workload consists of a mixture of 75% 4 KiB writes and 25% 16 KiB writes by setting the bssplit parameter in FIO, based on the observation that 25% of write requests in real-world cloud block storage workloads are at least 16 KiB and most of the write requests are smaller than 4 KiB [41]. We set the total write size of each workload as 64 GiB.
Figures 16(a)-16(d) show the throughput results. ZoneWrite-Only and ZoneAppend-Only cannot always achieve the highest throughput in all cases. For example, ZoneAppend-Only has 65.7% higher throughput than ZoneWrite-Only for 4 KiB writes and $(N_{s},N_{l})=(1,3)$ due to better intra-zone parallelism under small writes (Figure 16(a)), but ZoneWrite-Only has 27.2% higher throughput than ZoneAppend-Only for 16 KiB writes and $(N_{s},N_{l})=(1,3)$ (Figure 16(c)). On the other hand, ZapRAID achieves the highest throughput, or similar throughput to the best scheme, in all cases by carefully combining Zone Append and Zone Write.
Note that while ZoneAppend-Only can issue 16 KiB writes to multiple large-chunk segments, it cannot achieve higher throughput as $N_{l}$ increases, which may be inconsistent with the findings in Figure 22.1). One possible reason is that the computational overhead of the firmware for Zone Append becomes heavier and has higher fluctuations when we issue writes via Zone Append to multiple drives, and the performance is bottlenecked by the current slowest drive. Note that Zone Write also has a similar problem, but is less severe than Zone Append. Such performance fluctuations are also observed in prior work [75, 24].
Refer to caption Refer to caption
(a) 4 KiB writes, throughput (b) 8 KiB writes, throughput
Refer to caption Refer to caption
(c) 16 KiB writes, throughput (d) Hybrid 4 KiB and 16 KiB writes, throughput
Refer to caption Refer to caption
(e) 4 KiB writes, p95 latency (f) 8 KiB writes, p95 latency
Refer to caption Refer to caption
(g) 16 KiB writes, p95 latency (h) Hybrid 4 KiB and 16 KiB writes, p95 latency
Refer to caption Refer to caption
(i) 4 KiB writes, memory size (j) 8 KiB writes, memory size
Refer to caption Refer to caption
(k) 16 KiB writes, memory size (l) Hybrid 4 KiB and 16 KiB writes, memory size
Figure 16: Exp#11 (Multiple open segments). We show the throughput (figures (a)-(d)), 95th-percentile latencies (figures (e)-(h)), and stripe management memory usage per GiB of written data (figures (i)-(l)) under four open segments.
Figures 16(e)-16(h) show the 95th-percentile latency results. Overall, ZapRAID maintains low 95th-percentile latencies in all cases. For 4 KiB writes (Figure 16(e)) and hybrid 4 KiB and 16 KiB writes (Figure 16(h)), all three schemes have similar latencies except for $(N_{s},N_{l})=(1,3)$, ZoneWrite-Only has a significantly high latency since the 4 KiB writes are issued to the only small-chunk segment and ZoneWrite-Only has limited intra-zone parallelism.
For 8 KiB writes (Figure 16(f)), ZapRAID and ZoneWrite-Only have similar 95th-percentile latencies, while ZapRAID has 15.9-32.7% lower 95th-percentile latencies than ZoneAppend-Only for $(N_{s},N_{l})=$ (4,0), (3,1), and (2,2) by better exploiting inter-zone parallelism through both Zone Append and Zone Write. For $(N_{s},N_{l})=(1,3)$, ZapRAID has a slightly higher 95th-percentile latency than ZoneAppend-Only (by 9.4%) due to the overhead of group-based data layout.
For 16 KiB writes (Figure 16(g)), ZapRAID and ZoneWrite-Only have consistently lower 95th-percentile latencies than ZoneAppend-Only. We observe that the 95th-percentile latency of ZoneAppend-Only grows sharply by 173.3% from $(N_{s},N_{l})=(4,0)$ to $(N_{s},N_{l})=(3,1)$ since the 16 KiB writes are issued to the only open large-chunk segment for $N_{l}=1$ while they are issued to four open small-chunk segments for $N_{l}=0$. The 95th-percentile latency of issuing 16 KiB writes to a single large-chunk segment is much higher (see also Figure 9(c) in Exp#5). As $N_{l}$ further increases from one to four, the 95th-percentile latency of ZoneAppend-Only drops by 58.3% as there are more open large-chunk segments and a higher degree of inter-zone parallelism. We also observe that the 95th-percentile latencies of ZapRAID and ZoneWrite-Only increase by 121.1% and 115.3%, respectively from $(N_{s},N_{l})=(3,1)$ to $(N_{s},N_{l})=(2,2)$. The reason is that the Zone Write command on one open segment has low on-SSD computational overhead as it only handles one outstanding write request. The overhead grows when it issues multiple outstanding write requests to different open segments. Similar to ZoneAppend-Only, their 95th-percentile latencies drop as $N_{l}$ increases from two to four due to a higher degree of inter-zone parallelism.
Figures 16(i)-16(l) show the stripe management memory size per GiB of written data. We focus on ZoneAppend-Only and ZapRAID as in Exp#5. ZapRAID maintains significantly lower memory usage with its compact stripe table design than ZoneAppend-Only in all cases. For example, for 4 KiB writes (Figure 16(i)), ZapRAID achieves 68.9-98.5% lower memory usage than ZoneAppend-Only for $(N_{s},N_{l})=(4,0)$, $(3,1)$, $(2,2)$, and $(1,3)$. Also, ZapRAID incurs no memory usage for the compact stripe table for the cases of (i) $(N_{s},N_{l})=(0,4)$ and (ii) $N_{l}>0$ and there are only large writes (i.e., 16 KiB writes), since it only issues Zone Write to large-chunk segments. Thus, ZapRAID keeps low or no memory usage for stripe management, while ZoneAppend-Only always incurs significant memory usage for tracking the mappings between stripes and chunk locations.
[…]
Refer to caption Refer to caption
(a) $N_{l}=2$ and varying $N_{s}$, throughput (b) $N_{s}=2$ and varying $N_{l}$, throughput
Refer to caption Refer to caption
(c) $N_{l}=2$ and varying $N_{s}$, p95 latency (d) $N_{s}=2$ and varying $N_{l}$, p95 latency
Figure 17: Exp#11 (Multiple open segments). We compare the throughput and 95th-percentile latencies of RAIZN-SPDK and ZapRAID under hybrid 75% 4 KiB and 25% 16 KiB writes and different combinations of $(N_{s},N_{l})$.
We also compare ZapRAID with RAIZN-SPDK under multiple open segments. We focus on the workloads with hybrid 75% 4 KiB and 25% 16 KiB writes, and examine different combinations of $(N_{s},N_{l})$. We consider two settings: (i) we fix $N_{l}=2$ and vary $N_{s}$ from 0 to 6, and (ii) we fix $N_{s}=2$ and vary $N_{l}$ from 0 to 6.
Figure 17(a) shows the throughput results of RAIZN-SPDK and ZapRAID when we fix $N_{l}=2$ and vary $N_{s}$. For RAIZN-SPDK, when $(N_{s},N_{l})$ changes from $(0,2)$ to $(1,2)$, its throughput drops from 440.4 MiB/s to 268.6 MiB/s. The reason is that RAIZN-SPDK forwards all 4 KiB writes to two large-chunk segments for $(N_{s},N_{l})=(0,2)$, but for $(N_{s},N_{l})=(1,2)$, it forwards all 4 KiB writes to only one small-chunk segment and fails to leverage inter-zone parallelism for these small writes. When $(N_{s},N_{l})$ changes from $(1,2)$ to $(6,2)$, the throughput of RAIZN-SPDK increases to 1,077.6 MiB/s with the increasing inter-zone parallelism across small-chunk open segments. For ZapRAID, with the help of intra-zone parallelism for small writes, it maintains similar throughput for $(N_{s},N_{l})=(0,2)$ and $(N_{s},N_{l})=(1,2)$, even though the latter has only one small-chunk open segment.
Figure 17(b) shows the throughput results of RAIZN-SPDK and ZapRAID when we fix $N_{s}=2$ and vary $N_{l}$. The throughput of RAIZN-SPDK is 444.6 MiB/s for $(N_{s},N_{l})=(2,0)$, but only increases to 488.4-498.0 MiB/s when $(N_{s},N_{l})$ varies from $(2,1)$ to $(2,6)$. The reason is that most of the write requests (75% in total) are 4 KiB writes, and the performance is bottlenecked by the limited number of small-chunk open segments.
Figures 17(c) and 17(d) show the 95th-percentile latency results of RAIZN-SPDK and ZapRAID, and the observations are consistent with those in the throughput results. Overall, when there are multiple open segments, ZapRAID still achieves significantly higher throughput and lower latencies than RAIZN-SPDK, which incurs large overhead of partial parity updates.
($N_{s},N_{l}$) (0, 2) (1,2) (2,2) (6,2) (2,0) (2,1) (2,2) (2,6)
ZapRAID Wait phase 27.01 31.92 34.38 41.01 39.10 35.89 34.38 38.17
Data phase 72.86 74.30 66.86 67.25 76.31 66.23 66.86 70.27
Parity phase 5.44 3.69 3.68 3.52 3.43 3.76 3.68 3.60
RAIZN-SPDK Wait phase 775.46 1282.24 678.87 292.05 861.28 694.47 678.87 682.73
Data phase 23.66 22.62 23.96 31.46 26.08 23.83 23.96 24.24
Parity phase 1.01 1.02 1.02 1.11 1.04 1.04 1.02 1.04
Table 4: Exp#11 (Multiple open segments). Breakdown of latencies (in microseconds) of ZapRAID and RAIZN-SPDK for Figure 17.
Table 4 further shows the latency breakdown of RAIZN-SPDK and ZapRAID based on some of the settings of $(N_{s},N_{l})$ in Figure 17, so as to explain the large overhead of partial parity updates in RAIZN-SPDK. For each write request, we record four timestamps: (i) $t_{1}$, the time when the request is issued by fio, (ii) $t_{2}$, the time when the request starts writing data chunks, (iii) $t_{3}$, the time when the request finishes writing data chunks, and (iv) $t_{4}$, the time when the request completes writing all parity chunks (including partial parity updates for RAIZN-SPDK). We divide each write request into three phases: (i) the wait phase (i.e., $t_{2}-t_{1}$), (ii) the data phase (i.e., $t_{3}-t_{2}$), and (iii) the parity phase (i.e., $t_{4}-t_{3}$). The table shows that RAIZN-SPDK has a prolonged wait phase, as each of its write requests must update the parity chunks and wait for the completion of the partial parity updates of the previous write request for the same segment, thereby limiting the concurrency of writes. In contrast, ZapRAID has a short wait phase as it can issue writes concurrently (recall that our default queue depth is 64). Note that ZapRAID has longer data and parity phases than RAIZN-SPDK since its concurrent writes may overlap, so on average each write request takes longer to complete. Nevertheless, the overall write throughput of ZapRAID is significantly higher than that of RAIZN-SPDK.
Refer to caption Refer to caption Refer to caption
(a) 4 KiB chunks (b) 16 KiB chunks (c) Multiple open segments
Figure 18: Exp#12 (Overhead of garbage collection).
Exp#12 (Overhead of garbage collection in ZapRAID). We evaluate the impact of garbage collection on ZapRAID with different reserved space sizes. We issue writes to 200 GiB of data, with a total of 1 TiB of write traffic. We vary the physical space capacity of ZapRAID from 240 GiB to 400 GiB, meaning that the reserved space for garbage collection varies between 20% and 100%. We consider four write workloads, including 4 KiB (uniform) random writes, 64 KiB (uniform) random writes, 64 KiB skewed writes, and 64 KiB sequential writes. We consider 64 KiB writes to examine the overhead of garbage collection caused by much larger writes. For 64 KiB skewed writes, we consider a Zipfian distribution with a skewness factor of 0.99. We consider three settings: (i) a single open segment with 4 KiB chunks, (ii) a single open segment with 16 KiB chunks, and (iii) four open segments with $(N_{s},N_{l})=(2,2)$, $C_{s}=$ 8 KiB, and $C_{l}=$ 16 KiB.
Figure 18 shows the results. As expected, the write throughput increases (i.e., the garbage collection overhead is less) with more reserved space and with more skewed or sequential workloads. Since real-world workloads are often skewed in practice (e.g., block storage [72] and key-value [2, 8] workloads), we expect that the garbage collection overhead is acceptable in ZapRAID. Note that the write throughput under multiple open segments (Figure 18(c)) is higher than under single open segments (Figures 18(a) and 18(b)), meaning that the garbage collection overhead can be reduced via inter-zone parallelism.
Exp#13 (Overhead of L2P table offloading). We evaluate the overhead of offloading the L2P table entries to ZNS SSDs (§3.1). We consider the same four write workloads as in Exp#12 and focus on the setting of $(N_{s},N_{l})=(2,2)$ with four open segments. We set the physical space capacity as 300 GiB and issue writes to 200 GiB of data (i.e., the overprovisioned space is 50%). We vary the maximum memory size allocated for the in-memory L2P table from 50 MiB to 200 MiB. Recall that each L2P table entry has a size of four bytes for each 4 KiB block (§3.1). Thus, the in-memory L2P table can store the entries from 50 GiB to 200 GiB of data. Note that when the in-memory L2P table is allocated with 200 MiB, it can keep all L2P table entries in memory (i.e., no offloading).
Figure 19 shows the results. When the in-memory L2P table has less than 200 MiB of memory space, it needs to offload some entries onto the SSD. For 4 KiB and 64 KiB random writes, a 100 MiB L2P table has lower throughput than a 200 MiB L2P table by 59.2% and 16.8%, respectively. The reason for such high throughput drops is that under uniform random writes, the chance of not caching an L2P table entry is higher. Nevertheless, larger writes have lower degradations since ZapRAID groups the L2P table entries into mapping blocks and there will be less I/O overhead of reading L2P table entries from ZNS SSDs. For 64 KiB skewed writes, a 100 MiB L2P table only has slightly lower throughput than a 200 MiB L2P table by 4.0%, meaning that the offloading overhead becomes insignificant in skewed workloads. For 64 KiB sequential writes, a 100 MiB L2P table only has slightly lower throughput than a 200 MiB L2P table by 3.6%.
Figure 19: Exp#13 (Overhead of L2P table offloading).
5.5 Summary of Findings
ZapRAID is designed to integrate the strengths of both Zone Append and Zone Write primitives into a RAID system instead of always outperforming either in isolation. It is particularly effective for workloads with only small writes (e.g., 4 KiB or 8 KiB) or mixed workloads with most small writes (e.g., 4 KiB or 8 KiB) and few large writes (e.g., 16 KiB). This pattern is common in practice. For example, in the Alibaba Cloud traces (Exp#4), 684 out of 1,000 volumes (68.4%) contain more than 60% of requests no larger than 4 KiB and more than 25% of requests at least 16 KiB, with writes forming the majority of I/O.
Compared to ZoneWrite-Only, ZapRAID improves the write throughput for small writes in a single open segment, with a trade-off of introducing an extra compact stripe table. However, ZapRAID limits the size of the compact stripe table with the group-based data layout and ensures that the query overhead during degraded reads is minimal. While ZoneWrite-Only shows lower p95 latency in large writes in a single open segment (e.g., Figures 9(c) and 10(c) in Exp#5), ZapRAID introduces hybrid data management to use Zone Write for large writes and multiple open segments. When multiple open segments are available, ZapRAID shows no worse p95 write latency (Figure 16 in Exp#11).
Compared to ZoneAppend-Only, ZapRAID matches the write throughput in a single open segment because ZapRAID uses the Zone Append primitive for a single open segment. However, ZoneAppend-Only shows significantly poor degraded read performance (Exp#2) and higher memory usage for index management (Exp#3) in real-application experiments. ZapRAID outperforms ZoneAppend-Only with high degraded read performance (which is comparable to ZoneWrite-Only) and limited memory usage for index management.
We discuss the implications of our findings in terms of performance, memory usage, and functionalities of ZapRAID, compared with ZoneWrite-Only and ZoneAppend-Only.
Performance. ZapRAID demonstrates superior write performance and comparable normal read performance compared with ZoneWrite-Only and ZoneAppend-Only. Also, ZapRAID achieves significantly lower degraded read latencies than ZoneAppend-Only.
  • In terms of write performance, ZapRAID shows the highest write throughput under single-segment settings, as shown in Exp#1-Exp#4, Exp#5, and Exp#8, particularly with 4 KiB chunks, in which ZapRAID improves write throughput by up to 77.2% (Exp#5). Also, ZapRAID shows the highest write throughput and lowest write tail latencies across various configurations of $(N_{s},N_{l})$ under multiple-segment settings (Exp#1-Exp#4 and Exp#11). Furthermore, ZapRAID excels under highly concurrent write workloads (Exp#1 and Exp#10).
  • In terms of read performance, ZapRAID improves degraded read performance with lightweight stripe management. It maintains comparable degraded read performance compared with ZoneWrite-Only (Exp#6 and Exp#7) and has significantly higher degraded read performance than ZoneAppend-Only (Exp#2 and Exp#7). Note that ZapRAID is designed to maintain (rather than improve) normal read performance, so its overall throughput improvements are low for read-intensive applications (Exp#1 and Exp#4).
Memory usage. ZapRAID substantially reduces memory usage via its compact stripe table design compared with ZoneAppend-Only. It further reduces memory usage by offloading the L2P table to drives.
  • For the compact stripe table, ZapRAID reduces the compact stripe table size in two aspects. First, the group-based data layout (§3.2) significantly reduces the compact stripe table size for each zone by 66.7% (Exp#3 and Exp#5). Second, the hybrid data management (§3.3) leverages Zone Write for multiple-segment settings, so as to eliminate the need for the compact stripe table in those segments (Exp#3 and Exp#11).
  • For the L2P table, ZapRAID offloads the L2P table to drives to further reduce memory usage (Exp#13), with negligible overhead for highly skewed or sequential workloads.
Functionalities. ZapRAID supports multiple RAID schemes, crash recovery, full-drive recovery, and garbage collection, and hence adapts effectively to various scenarios.
  • ZapRAID maintains the highest write performance across various RAID schemes compared with ZoneWrite-Only and ZoneAppend-Only (Exp#8).
  • ZapRAID maintains high performance in crash recovery and full-drive recovery (Exp#9).
  • ZapRAID supports garbage collection by recycling overwritten or discarded user-written blocks. Its garbage collection overhead is small for highly skewed or sequential write workloads (Exp#12).
6 Related Work
SSD RAID. SSD RAID has been extensively studied in the literature. Some studies follow traditional disk-based RAID (with in-place updates) [52] and focus on reducing read tail latencies [19], improving parity update performance [21, 29, 13, 42, 24], optimizing garbage collection [33, 32, 71], making I/O performance predictable [40], and enhancing the scalability of Linux software RAID [69, 74]. To improve scalability, some studies organize RAID by distributing the controller function to multiple nodes [7] and further reducing the network traffic under disaggregated storage [61]. In this work, we focus on Log-RAID architectures, which have been extensively studied for SSD RAID to improve both write performance and flash endurance. SOFA [11] places the FTL in the RAID controller for efficient data management. Purity [14] manages both indexing and data storage under Log-RAID, and supports compression and deduplication. SALSA [23] implements a general translation layer for SSDs and Shingled Magnetic Recording (SMR) disks. SWAN [30] proposes spatial data separation to reduce garbage collection interference. Recent studies also apply RAID on Key-Value SSDs (KVSSDs) [54, 44, 56] and open-channel SSDs [43]. ZapRAID targets RAID for ZNS SSDs and particularly exploits Zone Append for high performance.
Storage systems for ZNS SSDs. Several studies explore new storage system designs for ZNS SSDs. Bjorling et al. [5] adapt F2FS [34] and RocksDB [58] for ZNS SSDs, and show that the adapted systems have lower I/O amplification and higher performance than conventional SSDs. Other studies focus on the performance and management aspects of ZNS SSDs. Examples include improving the performance of log-structured merge-tree stores [51] for ZNS SSDs [26, 37, 38], optimizing host-level garbage collection [12, 18, 60, 6], proposing new ZNS interfaces for efficient zone management [45, 18, 49], enabling ZNS SSDs for swap storage [4], designing new I/O scheduling for improved intra-zone parallelism [3], extending Zone Append for sub-block data appends [55], and ensuring crash consistency on F2FS backed by ZNS SSDs [35].
While the above studies focus on a single ZNS SSD, RAIZN [31] exposes a ZNS SSD array as a single ZNS interface to applications, and focuses on fault tolerance, correctness, and crash consistency. BIZA [73] presents a self-governing block-interface ZNS all-flash-array that hides the ZNS interface from applications, while internally exploiting zone random write area (ZRWA) and intra-/inter-zone parallelism to improve both endurance and performance. In contrast, ZapRAID focuses on exploiting Zone Append for high performance.
Interfaces for offloading address management. While our work focuses on the Zone Append interface, some studies propose new storage interfaces that offload address management from applications to storage devices as in Zone Append. Examples include range writes [1] and Direct File System [25] for file system management, nameless writes [76] and software-enabled flash [62] for flash-based SSDs, as well as ZEA [46] for host-managed shingled magnetic recording drives. We conjecture that ZapRAID, which builds on Zone Append, can also build on the above interfaces to improve RAID performance, and we pose this issue as future work.
7 Conclusion
ZapRAID is a high-performance Log-RAID system for ZNS SSDs to support scalable and reliable storage, with the design goals of high performance, lightweight stripe management, and reliability in mind. We propose the following key design elements: (i) a group-based data layout under Zone Append to effectively exploit intra-zone parallelism while mitigating the stripe management overhead; (ii) hybrid data management to carefully combine Zone Append and Zone Write to achieve intra-zone and inter-zone parallelism; (iii) the crash consistency mechanisms to enforce high reliability. Our prototype evaluation on real ZNS SSDs and FEMU emulation shows that ZapRAID achieves high write throughput and maintains efficiency in degraded reads, crash recovery, and full-drive recovery.